### v1

# Upgrading Democracies with Fairer Voting Methods

Abstract
11footnotetext: Corresponding author: Evangelos Pournaras, School of Computing, University of Leeds, Leeds, UK, E-mail: e.pournaras@leeds.ac.uk
Voting methods are instrumental design element of democracies. Citizens use them to express and aggregate their preferences to reach a collective decision. However, voting outcomes can be as sensitive to voting rules as they are to people’s voting choices. Despite the significance and inter-disciplinary scientific progress on voting methods, several democracies keep relying on outdated voting methods that do not fit modern, pluralistic societies well, while lacking social innovation. Here, we demonstrate how one can upgrade real-world democracies, namely by using alternative preferential voting methods such as cumulative voting and the method of equal shares designed for a proportional representation of voters’ preferences. By rigorously assessing a new participatory budgeting approach applied in the city of Aarau, Switzerland, we unravel the striking voting outcomes of fair voting methods: more winning projects with the same budget and broader geographic and preference representation of citizens by the elected projects, in particular for voters who used to be under-represented, while promoting novel project ideas. We provide profound causal evidence showing that citizens prefer proportional voting methods, which possess strong legitimacy without the need of very technical specialized explanations. We also reveal strong underlying democratic values exhibited by citizens who support fair voting methods such as altruism and compromise. These findings come with a global momentum to unleash a new and long-awaited participation blueprint of how to upgrade democracies.
1 Introduction
Reaching collective decisions that benefit all often remains an unfulfilled promise. While voting methods have been instrumental to adaptively govern societies, they are often based on outdated approaches to foster collective intelligence [1, 2, 3]. This may put democracies at risk instead of empowering them [4]. Over-relying on limited-choice majoritarian voting often results in promoting competition instead of cooperation, in marginalization of minorities instead of inclusion, in few winners and many losers [5, 4, 6, 7, 2, 3, 8]. Polarization, exacerbation of inequalities, and radicalization of the public sphere [9, 10] are prominent implications of voting methods that fail to foster fairness by design.
The challenge of adopting democratic innovations. Proportional voting methods and collective decision making have attracted broad interest by inter-disciplinary research communities [11, 12, 13]. This includes the fields of social choice, computational social science, economics, complexity theory and political science. Nevertheless, adopting the developed methods in practice and changing the status quo is itself a long-standing and multi-faceted challenge [14]. It often comes as a ‘chicken and egg’ problem: difficulties in communicating alternative voting methods effectively and lack of evidence of how they function in real world cause risk aversion at the political level and limited will for change, which further maintains the lack of evidence, also in regards to the legitimacy perception of voters [15, 16, 17, 2, 10]. Such unprecedented complexity for democratic innovations requires rigorous scientific methods [17, 14, 10, 18]. A blueprint for shifting to fairer voting methods is urgently needed.
Participatory budgeting as a blueprint for democratic innovations. In recent years, there has been quite a number of noteworthy “democratic innovations". These include citizen assemblies, government labs (“GovLabs"), open innovation (open access/data/source, creative commons), crowd sourcing/sensing/funding, hackathons, fablabs, makerspaces, and citizen science. While related to some of these recent trends, our article focuses on significant interventions in participatory budgeting. This is a process designed to distribute a public budget of a city council in a bottom-up way by engaging citizens in collaborative brainstorming sessions to propose their own projects, solutions, and ideas, on which they later vote [19]. Citizens may also be involved in the implementation of selected projects. Participatory budgeting offers a great potential for demonstrating grass-roots democratic innovations in real-world, which can be the blueprint to scale up further in other democratic processes. Started in Porto Alegre in 1989, participatory budgeting has since been adopted in many places across the globe, from remote communities in Kenya, Indonesia, and the Philippines to major cities such as Barcelona, Warsaw, Helsinki, Paris, and New York [20, 21]. Its uptake has been particularly significant in Europe, with a surge in participatory initiatives in recent years [19]. By 2019, Europe accounted for more than half of all participatory budgeting schemes worldwide, with more than 5,000 initiatives [21]. Despite this uptake, the impact of fairer voting methods on voting outcomes as well as project ideation and implementation are, so far, under-explored questions. This is despite their profound importance for the quality of the voting outcomes.
Collective choice made fairer. We demonstrate a breakthrough for democracies by applying two alternative voting methods designed for a more expressive and proportional representation of citizens in a real-world setting. This has been achieved with a novel field test designed for a systematic assessment of a new participatory budgeting process in the city of Aarau in Switzerland. We used the ballot format of cumulative voting [22, 23] allowing voters to distribute 10 points over at least 3 out of a total of 33 project ideas. Compared to other prominent input voting methods such as approval, ranking, or knapsack voting [24], cumulative voting enables voters to express their preferences more effectively by weighting their support to different options. This expression of preferences is more inclusive and can provide a more proportional representation of minority groups. Complementary, we used the voting method of equal shares [25] to aggregate the points allocated by citizens and elect the winning project ideas. Equal shares promotes a more proportional representation in comparison with majoritarian voting. This is favorable, as majoritarian approaches to elect winners can sometimes lead to a "tyranny of the majority", where "the winner takes it all", i.e. monopolizes all the budget [26]. Proportional representation is achieved by assigning to every voter a decision power, which is given by an equal share of the budget. This manifests both in meeting voters’ preferences and in the geographical coverage of the winners [2]. In practice, the method of equal shares is resourceful, as it is likely to select a larger number of projects, which are cost-efficient, while satisfying the preferences of more people and increasing overall voter satisfaction [27, 28]. Given the applied principle of proportionality, we refer to these alternative decision-making approaches here as ‘fair voting methods’. As it turns out, they also promote legitimacy and inclusion. Our study provides universal findings, beyond Aarau, as we also study a large number of past participatory budgeting processes and recent follow ups adopting fair voting methods.
Overcoming barriers of shifting to fair voting methods. Combining cumulative voting and equal shares in a real-world participatory budgeting process allows us to unravel: (i) how to systematically adopt fair voting methods. Apparently, for policy-makers and a general public familiar with elections, the method of equal shares is pretty different from the widespread and longstanding standard method of utilitarian "greedy" voting, which simply sums up the votes and selects the most popular project ideas fitting the budget in a majoritarian way; (ii) how the method of equal shares influences participation in the project ideation phase, allowing for the consideration of additional criteria during the voting process. Project costs become instrumental, as more expensive projects require broader support. This may discourage expensive projects, which could be compensated by a larger number of more resource-efficient projects which are fairer distributed in space [27]; (iii) the legitimacy of the voting outcomes. Although the method of equal shares is designed to promote fairness, sacrificing popular expensive projects may be perceived as unfair by their supporters, who may be a considerable share of the population. Whether the selection of more alternative projects can compensate for the potential disappointment is key for the perceived legitimacy of fair voting methods; (iv) causal factors that promote a shift to fair voting methods. The crucial question is, whether there are any personal traits or societal values that support and accelerate democratic upgrades based on fair voting methods.
A systematic and rigorous approach to upgrade democracies with fair voting methods. We co-designed the participatory budgeting process of "Stadtidee" ("City Idea") in collaboration with the city of Aarau in Switzerland. It was conducted in 2023 with an ongoing project implementation. It is characterized by the following innovations (see Section 4.1 for more information): (i) the real-world application of the method of equal shares in combination with cumulative voting as the ballot format. Taken together, they enable a better, proportional representation of citizens’ preferences; (ii) an inclusive participation fostered throughout the entire process. The citizens proposed 161 ideas, out of which 33 were voted on by 1703 citizens, women in majority and children included, with 17 projects selected for implementation [29], see Table 1 for an outline; (iii) a survey study was conducted before and after voting. This allows us to rigorously assess the new process in depth over its entire lifetime. We are also able to assess the legitimacy of fair voting methods before and after voting and to profoundly understand the dynamics of voting outcomes that these methods generate, including causal and explanatory factors; (iv) a secure and privacy-preserving voting using and extending the Stanford open-source platform for participatory budgeting [30, 31]. This ensures transparency and trust when using fair voting methods. These four aforementioned innovations offer a strong foundation of validity to rigorously test fair voting methods. They provide a blueprint for testing upgrades of democracy, which have been missing so far.
2 Results
This articles presents three key results:
1. Fair ballot aggregation results in voting outcomes with more winners, while achieving better representation of voters’ preferences and novel project ideas, also in space.
2. Voters find the adopted ballot aggregation method of equal shares more legitimate, by preferring it over the majoritarian standard method and finding it fairer.
3. The shift to a fairer voting method distinguishes supporting voters with strong democratic values from voters without these values. Democratic values that explain the preference for more proportional voting methods include altruism, opposition to costly popular projects, and support for more expressive and inclusive ballot formats.
2.1 Fair ballot aggregation has distinguishing multi-faceted qualities
Why the method of equal shares results in more winners. Figure 1a shows the 33 projects proposed for voting and the related voting results. The following key observations can be made: (i) Equal shares selects 17 winning projects, with more than half of the proposed projects winning. In contrast, the standard utilitarian greedy method would select 7 projects only, which would imply 26 losing projects, thereby creating a significant deficit in representing the broad range of citizen preferences. (ii) The project entitled ‘Wild bees’ paradise’ is an example of an expensive popular loser using the method of equal shares, which would win under the standard method. With 33.64% of voters selecting this project, it is the 3rd most popular project, which would consume 40% of the overall available budget. The method of equal shares selects several other, lower-cost projects instead, each of which may be less popular, but they address together the preferences of a broader population. (iii) The mean budget share of the proposed projects is 14.33%, which is 48.37% lower than a typical proposed project within 345 past participatory budgeting elections conducted with the standard method [32]. This illustrates the impact of ballot aggregation design already in the project ideation phase: it is more resourceful by encouraging lower-cost projects of different and often novel nature, with the potential for synergetic impact [27].
Figure 1: Compared to the standard utilitarian greedy method, which is majoritarian in nature, the number of winning projects resulting from the method of equal shares increases dramatically, with more than half of the proposed projects winning, while citizens’ proposed project ideas are better represented geographically. (A) Equal shares elects 17 winning projects compared to 7 with the majoritarian standard method. The project ‘Wild bees’ paradise’, requiring 40% of the total budget, is popular winner under the standard method but loser under equal shares. This allows more low-cost projects to win, which, in combination, better represent voters’ preferences. (B) Under equal shares, five districts receive more winning projects as compared to the standard method, thereby strengthening the periphery of the city. (C) Spatial fairness increases by 44% under equal shares as measured by the relative reduction in the index of dispersion for the number of winning projects relative to the proposed ones for each district.
Equal shares improves voters’ geographic representation, especially for city peripheries. Figure 1b illustrates that the method of equal shares results in 5 districts (27.77%) with more winning projects compared to the majoritarian standard method, while no district receives a lower number of projects. This particularly benefits the periphery of the city that would otherwise be under-represented. Moreover, Figure 1c measures the relative number of winning/proposed projects in each district. Using the method of equal shares, the spatial fairness increases by 44% when measured by the reduction in the index of dispersion, which demonstrates that more proposed project win in each district. These results confirm that the periphery of Aarau is better represented, both in terms of the voter preferences and the number of proposed project ideas selected.
The method of equal shares represents voters better. Figure 2a shows the share of voters for different cumulative levels of the mean representation of their preferences. Representation is measured by the relative number of projects that are preferred by a voter and win. Comparing the standard method of 345 past participatory budgeting processes [33, 32] with the outcomes for the equal shares method in Aarau, we observe 3.7% more voters with higher representation levels, in particular for lower levels of cumulative representation. This means that equal shares better preserves the representation of the population even for under-represented voters, whose preferences do not align well with the voting outcomes. Specifically, for the participatory budgeting process in Aarau, the voters’ representation is 19% higher under equal shares compared to the standard method and 14.5% higher on average compared to 345 earlier participatory budgeting processes.
Figure 2: Compared to earlier participatory budgeting experience based on the standard utilitarian greedy method, the method of equal shares increases the representation of voters’ preferences and project categories, in both project ideation and selection phase. (A) The representation of voter preferences increases by 14.5% under equal shares compared to earlier participatory budgeting processes based on the standard utilitarian greedy method. (B) The projects proposed for voting under equal shares promote project categories that have traditionally been under-represented, such as welfare and health, while projects that usually come with high costs, such as public transit, lose appeal significantly. (C) It is also these otherwise under-represented projects that become the winning ones when using equal shares. On average, project categories show a 3% higher representation.
Equal shares encourages proposed projects that are novel and win. Figure 2b illustrates the share of the proposed project categories of participatory budgeting in Aarau compared to the projects observed in earlier participatory budgeting processes [33, 32]. We find a significant change in the nature of the proposed projects: welfare and health projects, which were usually under-represented, are more prominently featured, while public transit projects are less prominent. This shift is determined by the costs of the projects belonging in the two categories: the average cost of public transit projects in the participatory budgeting process of Aarau amounts to 8% of the overall budget, while welfare-related projects amount to 12.25%. In earlier participatory budgeting processes, these numbers were 18.19% and 7.52%. Figure 2c presents the share of the winning project categories in the participatory budgeting process of Aarau for the method of equal shares and for the majoritarian standard method. It also shows the share in earlier participatory budgeting experience that relied on the standard method. Health and sport would be disadvantaged with the standard method as it would reduce their share in the winning set by 19.13% and 18.05%, respectively.
Broader evidence beyond the case of Aarau. To strengthen further our findings, we also highlight the case of two other participatory budgeting elections that used equal shares during 2023 in two cities of Poland: (i) Świecie and (ii) Wieliczka (“"Green Million"). In the first case that was a follow up of Aarau, 17/22 projects are selected for funding, while the majoritarian standard method would result in only 9 winning projects. The representation of voters’ preferences increased from 61% to 80%,. The top-2 project in voters’ approvals is not funded, which requires 36.6% of the budget and is approved by 24.4% of the population. In the second case, 30/64 projects are selected for funding, while the majoritarian standard method would result in 23 winning projects. The representation of voters preferences increased from 61% to 71%,. The top-2 project in voters’ approvals is not funded, which requires 10% of the budget and is approved by 8.9% of the population. Both cases confirm the distinguishing voting outcomes of equal shares: proportional representation is achieved by funding more projects that increase the overall satisfaction with compromises on popular expensive projects. In contrast to these two elections, further novel data are collected during the participatory budgeting process of Aarau that allow us to rigorously explore the legitimacy of fair voting methods and which citizens’ values explain the shift to those.
2.2 Overcoming the adoption barrier of fair voting improves legitimacy
Adopting a new ballot aggregation method: a rigorous evaluation. Figure 3 illustrates which ballot aggregation method voters prefer and which they find fairer in the entry survey before voting and in the exit survey after voting. Section LABEL:sec:choice-ballot-aggregation also shows additional results of which method voters find more convenient. Section LABEL:sec:equal-shares-explanation further illustrates how the two methods are presented to voters, in particular the voting outcomes and the explanations, based on which the measurements are made. The shift from the majoritarian standard method to the method of equal shares becomes obvious by the gain in the share of voters, who prefer and find fairer equal shares after voting (rigorously measured on the same matched population). We also assessed which method voters found more convenient in Figure LABEL:fig:convenience. The legitimacy of the two ballot aggregation methods is determined by measuring the satisfaction of voters with the voting outcomes of each method (even if the voter classifies as a ‘loser’, based on different measurements), see Section 4.2 for more information. To assess legitimacy thoroughly, the voters’ gain is disaggregated further down to all 24 combinations of winners/losers and satisfied/unsatisfied voters (see Figure LABEL:fig:legitimacy_fairness_preference_heat). We group those into three groups in Figure 3 for simplicity: satisfied winners, unsatisfied losers and mixed, each consisting of 8 combinations. Three ways to measure each of the winners/losers and satisfied/unsatisfied are illustrated in Section 4.2.
Figure 3: In the exit survey after voting, the method of equal shares is found more preferred and fairer than the majoritarian standard ballot aggregation method. Strikingly, before voting and adopting equal shares, there is no clear preference for any of the two methods, while the standard method is found fairer. Voting outcomes themselves are sufficient for individuals to determine equal shares as their preferred and fair ballot aggregation method. The explanation of the methods does not significantly contribute to shifting the choice to equal shares. The legitimacy of equal shares is assessed based on voters’ (A) preference and perceived (B) fairness of equal shares over the standard method. The x-axis determines the difference of the voters share between those choosing equal shares over the standard method. Legitimacy is assessed based on 24 combinations of winners/losers and satisfied/unsatisfied voters that are grouped into (i) winners and satisfied, (ii) losers and unsatisfied and (iii) mixed. Winners/losers are determined in three different ways: 1st and 2nd quartile of representation score for the winning projects that are (i) non-exclusive and (ii) exclusive between the two ballot aggregation methods. (iii) The non-supporters and supporters of the top popular losing project. Satisfied/unsatisfied voters are determined by three post-voting survey questions that assess: (i) satisfaction, (ii) acceptance of outcome and (iii) accurate representation of will.
The adoption of equal shares exhibits a shift to be the most preferred and fairer method. The results of Figure 3 show that, in the exit survey after voting as compared to the entry survey before, there is a significant gain in voters who prefer the method of equal share and find it fairer. This is particularly prominent for unsatisfied losers who choose the standard method as fairer before voting but switch to equal shares after voting (see Figure LABEL:fig:legitimacy_fairness_preference_heat). More specifically, 54% of the population prefers equal shares before voting, while after voting, supporters of equal shares raise to 80% (p=0.0034). Similarly, 45% of voters find equal shares fairer before voting and 89% fairer after voting (p=0.028).
The method of equal shares shows stronger legitimacy. The significant legitimacy of equal shares is reflected in several ways (see also Figure LABEL:fig:legitimacy_fairness_preference_heat): (i) Losers, determined as supporters of the top popular losing project (‘Wild bees’ paradise’), prefer the method of equal shares and find it fairer. (ii) Remarkably, the method of equal shares is preferred and found fairer even more among losers than among winning voters, who prefer the majoritarian standard method. (iii) The shift to equal shares after voting is even found among those who do not accept the voting outcome, which is a strong proxy of legitimacy [34, 35].
When shift to equal shares is not achieved. However, we also found some cases where more voters shifted to the majoritarian standard method after voting, namely (i) satisfied voters preferring the standard method over unsatisfied voters preferring equal shares; (ii) voters accepting the voting outcome and finding the standard method fairer over voters who do not accept the voting outcome and find equal shares fairer; (iii) unsatisfied losers (or winners), who prefer the standard method over unsatisfied winners (or losers), who choose equal shares, when winners/losers are determined by the representation and their choice for the top popular losing project (‘Wild bees’ paradise’). See Figure LABEL:fig:legitimacy_fairness_preference_heat for a more disaggregated view.
Voting outcome itself vs. method explanation: what matters. The method explanation has an additional contribution of 5.3% to the shift to equal shares as preferred method, while it has an addition contribution of 5.2% to the shift to equal share as fairer method. The method explanation has a small but statistically significant contribution to the shift to equal shares (with only exception the mixed setting for the preferred method, p=0.07). These results demonstrate that the voting outcome itself is sufficient to establish a change.
2.3 Democratic values explain voters’ shift to fairer voting methods
Towards fairer voting: A clash between democratic values and democratic backsliding. Figure 4 identifies distinguishing human factors that causally explain the shift from one ballot aggregation method before voting to another one after voting (with the explanation of method), as illustrated in Figure 3 and in Section 4.3. The classification model is illustrated in Section LABEL:sec:shifts-methods and comes with an accuracy of 85.92%. Strikingly, the group of voters who shift from the majoritarian standard method to equal shares exhibits democratic values that explain this transition, see Figure 4c. These voters come with distinguishing qualities (Figure 4b) such as being altruistic, compromisers and supporting more expressive, inclusive ballots. They are aware of the proposed projects and flexible to select low-cost projects, which also makes them more likely to oppose a popular costly project. This group is well-represented in the voting outcome. In contrast, the voters that shift from equal shares to the majoritarian standard method do not exhibit these democratic values, but are driven by self-interest, mistrust and dissatisfaction. They support simpler ballot formats, are not well-represented in the voting outcomes, and support the popular loser project under equal shares.
Figure 4: The shift to equal shares as the preferred method after voting is explained by distinguishing human traits that reflect the promotion of strong democratic values such as being altruistic, compromisers and supporting more expressive, inclusive ballots. In contrast, shifting to the majoritarian standard method is found among voters showing mistrust, under-representation, dissatisfaction, and self-interest. These voters also support the popular costly loser project of the equal shares and simple ballot methods. A classification model (with 85.92% accuracy) is designed with four classes representing the preferred ballot aggregation method before and after voting (with explanation of methods). The independent variables are human traits of voters that explain the observed transitions. (A) Distinguishing human traits of voters who shift to the standard method and equal shares after voting. These are prominently associated with democratic backsliding on the one hand, and promotion of democratic values on the other hand. (B) Summary of distinguishing qualities for those who prefer the method of equal shares over the majoritarian standard method after voting (feature difference of the two classes). Altruism, support of expressive, inclusive ballots and opposition to the popular, costly loser project of equal shares are qualities with the highest prominence for those voters preferring equal shares. (C) Human traits (features) that explain the preference of each voter before and after voting, in particular the transitions from the majoritarian standard method to equal shares and from equal shares to the standard method (see Figure 3). The ‘*’ indicates that a feature is statistically significant (p<0.05) in explaining the voters’ choice of the method before and after voting.
Distinguishing qualities of voters who shift preference to equal share. Figure 4b summarizes the distinguishing qualities of voters, who shift to equal shares by taking the difference of the human traits for the two transitions (standard to equal shares - equal shares to standard). Altruism, opposition to the popular (costly) loser project and support of the expressive, inclusive ballot formats are the most prominent distinguishing qualities observed for voters who shift to equal shares. Figure 4c shows the importance of features that explain the classification of voters based on which of the two transitions they make for their preferred ballot aggregation method before and after voting (see Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-loss and LABEL:tab:trait-survey-map in Section LABEL:sec:shifts-methods). Further results for the transitions to methods that are found fairer and more complex are shown in Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
3 Discussion
A blueprint to upgrade democracies. In our study, we have shown how voting can better foster collective intelligence that is based on higher inclusion and diversity, with strong evidence provided by a rigorous participatory budgeting process in the city of Aarau in Switzerland. Subsequent participatory budgeting processes run using this approach further confirm our findings [36]. In particular, we found that the majoritarian standard approach of voting had a number of deficits, such as the selection of a few projects, leaving a large portion of the project ideas unfulfilled. As the selected projects are implemented mainly in the city center, it largely fails to reach spatial fairness or inclusion. However, combining cumulative voting with the method of equal shares proves to be resourceful, as it promotes more novel winning projects, fairness, satisfaction, inclusion, trust and democratic values. Accordingly, this approach offers a powerful way of upgrading democracies. Implementing this approach, using digital means, promises to be a powerful, value-sensitive approach to promote digital democracy by design [7].
Success of project implementation. By the end of 2024, 9 of 17 projects selected by the "City Idea" process in Aarau were already implemented and 5 projects got an extension, while 3 projects could not be implemented mainly due to feasibility and budgetary constraints [29]. These were among the top-11 projects. Two thirds of these projects would also have been winners if the majoritarian standard method had been used. These results demonstrate that having more winning projects using the equal shares method does not negatively affect the overall capacity of citizens to implement them.
Participation levels. The participatory budgeting process in Aarau achieved significant participation of citizens at different levels. During the project ideation phase, 161 ideas were proposed, which exceeded the initial goal of the campaign to attract 100 ideas by far [29]. The total of 1703 voters who participated is a significant number for such a campaign in a town with 14,310 residents having voting rights in 2023. Nevertheless, the turnout did not reach the level of official elections such as referenda and national elections. A 22% of the voters showed interest in participating, although they were not officially residents of Aarau, while 9% of the voters were not Swiss citizens. The average age of voters was 43, and there were more female than male voters. Children over 15 years old were allowed to vote. Among the low number of voters who abstained but participated in the surveys (296), the top-5 reasons for not voting were "no time" (17.1%), "forgetting" (12.8%), "lack of information" (9.9%), "unappealing projects" (7.7%) or some "other reason" (7.0%), see Question EB.7 in Table LABEL:table:exit-digital.
Upgrading democracies requires upgrading digital tools. We created and used a novel combination of open-source and value-sensitive digital tools to meet a broad range of expectations regarding transparency, security, privacy, reproducibility, explainability and fairness. This required a significant implementation effort, a mobilization of various communities, a strong commitment and the building of shared values with the city authorities, which is a process that has not been systematic pursued to this extent so far. A highlight of this effort was the privacy-preserving verification of citizens’ eligibility when voting with the social security number, which ensured a trusted voting procedure. The integration of our open-source implementation of cumulative voting [30] into the Stanford Participatory Budgeting platform is a stepping stone towards a more ambitious roadmap, which aims to make participatory tools more standardized and interoperable, in particular the ones supported by large open-source communities such as Decidim and Consul as well as the ones created by the computational social choice community such as Pabutools [37] and Preflib [38].
Explaining and adopting fair voting methods. A 53.17% of voters understood how equal shares calculates the winning projects (MB.3.3), while a 55.95% understood why popular projects are not easily funded by equal shares (MB.3.6), and 41.26% find that fair (MB.3.7). Strikingly, the results of Figure 3 demonstrate that explaining the expected difference in the voting outcomes reached by equal shares is sufficient for voters to prefer this method. A more detailed explanation of how the method works increases the tendency towards preferring equal shares over the standard method only slightly. This result has significant implications for overcoming the long-standing hesitation with regard to democratic innovations, since citizens show a readiness to accept more complex decision-making methods as long as they understand the anticipated impact on the decision outcome. In addition, it helps when people experience the different result, which gives them a practical understanding of the method. Moreover, given that equal shares is designed for a proportional representation of voter preferences and is known for its stability [19], it is more unlikely to yield controversies over voting outcomes. Nevertheless, the additional effort required to explain the method of equal shares should not be under-estimated. As shown in Section LABEL:sec:equal-shares-explanation, the introduction of equal shares requires a systematic communication effort over the lifetime of the participatory budgeting process, using creative media and visualizations suited for the broader public. At the project ideation phase, awareness of the method is required to mobilize different parts of the populations for ideas that may not have been considered under the majoritarian rule. As the costs of the proposed projects play a key role for selecting them under proportional representation, citizens need to be resourceful and prioritize different ideas, meaning a focus on cost-effective projects that get the support by a sufficiently large group of the population. Large and costly projects may be more risky to propose as they require the support of a broader population to get selected.
Fair voting methods to foster a novel impact creation process. The influence of the chosen voting method on the proposed and selected project ideas can transform the process of creating real-world impact. For instance, novel welfare projects gain momentum, as they appeal to several interest groups, while they may also be easier to implement than costly infrastructural projects [27]. We foresee that, on the long term, proportional voting methods could enable an incremental grass-roots implementation of sustainable development projects over several participatory budgeting campaigns, which could also support a fairer transition to a net zero future.
Fair voting methods: a better-than-optimal paradigm. Over-simplifying, averaging or overriding people’s preferences while optimizing societies for goals that people may not even explicitly agree on is a significant risk that democracies face in the era of artificial intelligence (AI) run by Big Tech. In this context, the notion of ‘optimal’ solutions is actually misleading and displaced from the reality of people, as demonstrated by the inferiority of popular expensive projects or projects focused on city centers that are typically proposed in participatory budgeting processes under majoritarian standard voting. In contrast, proportional voting methods work well for pluralistic settings, where voters may not agree on one and the same goal, but decisions are based on collective intelligence. This approach excels when goals are complementary, but is also suited to handle situations of competitive or even incompatible goals, or when significant minorities are involved. Small compromises, such as excluding a popular expensive project, open up opportunities for large collective returns with benefits for a great majority of people. For instance, the inclusion of under-represented health projects or of projects in the city periphery can reduce social tensions and benefit a broader population. The demonstrated approach of combining cumulative voting with the method of equal shares delivers a powerful blueprint for upgrading democracies. It allows societies to achieve (i) a larger number of realized projects within the respectively available budget; (ii) a greater resourcefulness of novel projects and a multi-faceted impact; (iii) more participation and inclusion; (iv) greater transparency and trust in the voting process; (v) better representation of citizen preferences and higher satisfaction; (vi) more (spatial) fairness; (vii) higher legitimacy thanks to a greater acceptance of the collective outcomes; (viii) strengthened democratic values of voters. We refer these impacts "better than optimal" as they tend to be a win-win for all, while they are distinguished from what existing optimization and data-driven decision-making approaches typically offer.
Fair voting methods in the era of artificial intelligence. We would like to end on a remark addressing the use of data-driven AI systems for decision making. While we consider AI a very powerful technology with numerous promising application areas, how to use it in the context of governing democracies must be studied very carefully. AI safeguards for democracy are critical and yet to be created. Existing decision-making methods supported by AI solutions show alarming vulnerabilities and are still perceived as lacking fairness and legitimacy [39, 40, 41]. For instance, biases and inconsistencies by unconsolidated use of large language models can exacerbate by voting methods with a deficit on fairness [42]. It is not clear when and to what extent AI risks can be addressed. Therefore, alternative approaches such as the one presented here are urgently needed.
4 Methods
We illustrate the designed participatory budgeting campaign of City Idea, the evaluation approach of legitimacy and the models for causal explanations for the shifts between ballot aggregation methods before and after voting.
4.1 The participatory budgeting campaign of City Idea: field test design
Table 1 outlines the participatory budgeting campaign of City Idea. It consists of six phases over the passage of nine months. An entry and exit survey were conducted before and after voting respectively. The entry survey was launched via physical invitation letters to all citizens sent by city council (10,000 A5 flyers [29]). This resulted in 3592 respondents. Some of these filled in their email to voluntarily participate in the exit survey and this resulted in 808 respondents. The surveys included questions about demographics and political opinion (Table LABEL:table:entry-demographics-LABEL:table:exit-demographics), engagement and digital skills (Table LABEL:table:entry-digital-LABEL:table:exit-digital), the City Idea voting (Table LABEL:table:entry-city_idea-LABEL:table:exit-city_idea), the voting outcome preference (Table LABEL:table:entry-voting_outcome-LABEL:table:exit-voting_outcome) and the voting input and aggregation method (Table LABEL:table:entry-voting_input-LABEL:table:exit-voting_input). Survey questions encode variables, which are observed for changes before and after voting. The overall study has received ethical approval from University of Fribourg (#2021-680).
Table 1: The rigorously designed participatory budgeting campaign of City Idea with 6 phases over a period of 9 months in 2023. Participants of the entry survey, voting and exit survey are linked using an anonymous identifier (hash) of the citizens’ social security number. This allows the study of several choice variables before and after voting.
Stage Start End Scale Shared Participants in Stages
1. Entry survey January March 3592 participants
2. Project ideation 13th of February 18th of March 161 projects
3. Deliberation 20th of March 11th of June 33 projects
4. Projects campaigning 5th of June 11th of June 73 participants, 1682 comparisons
5. Voting (All) 12th of June 25th of June 1703 participants
5. Voting (Aarau) 1330 participants
6. Exit survey July September 808 participants
Matched population: 505 505 804 252 252 252 252
Implementation July 2023 December 2024 9 implemented, 5 extended
Study of ballot formats. The preference over different ballot formats is studied via Question MB.2 in exit survey (see Table LABEL:table:exit-voting_input). The transitivity over different ballot formats is studied with a hypothetical voting scenario in the entry survey with 5 costed projects (Question MA.6 to MA.10 in entry survey) and a budget of 50,000 CHF, in which participants voted using 5 different ballot formats in sequence, see Figure LABEL:fig:ballot-formats. The involved ballot formats include: single choice, approval, score, Knapsack, and ranking.
Study of ballot aggregation. The preference, fairness and simplicity of the standard method versus the method of equal shares are assessed (Questions MA.1, MA.3, MA.4 in entry survey and Question MB.4, MB.6, MB.7, MB.9, MB.10 in exit survey). Before voting, participants assess the voting outcome of a simple hypothetical voting scenario depicted in Figure LABEL:fig:ballot-aggregation. Post voting, participants assess the legitimacy of aggregation methods on the real-world ballots after exposure to the winning projects elected via ballot aggregation methods followed by the explanation of the methods (see Figure LABEL:fig:q2224graphic). With this two-stage assessment we can distinguish the choice of an aggregation method because of its voting outcomes vs. the method explanation.
Project ideation, deliberation and campaigning. Citizens proposed more than 161 project ideas out of which 33 projects are selected to put for voting [29]. Projects are eliminated or similar project ideas are combined to strengthen their chances to win. Project ideation involved support and guidance with planning, implementation, economic and legal feasibility by a municipal advisory board of 7 people, a core team of 3 people and ambassadors. The city council distributed posters and flyers to various municipal institutions and locations. It also organized two press releases that resulted in 17 media reports. Moreover, the interactive information campaign ‘know your projects’ was launched in the city that engaged 73 participants in 1682 random pairwise comparisons of projects (see Table LABEL:tab:pairwise_winners for further information) via the Stanford Participatory Budgeting platform  [43]. Awareness about the method of equal shares and what projects are expected to win was raised, emphasizing the challenge of electing projects that cover a large share of the budget. Section LABEL:sec:equal-shares-explanation provides more information about how the method of equal shares was communicated and explained to citizens.
Voting. The city of Aarau has a total of 22,593 citizens, out of which 17,428 are Swiss citizens. A total of 1703 voters participated in the City Idea campaign, out of which 1330 are citizens of Aarau. This also allows individuals to have a voice, who are not living in Aarau, but have a connection to the city, such as workers and broader family members. Children over 15 years old could also vote. Voting was digital, using the open-source Stanford Participatory Budgeting platform [43]. However, voters with low digital literacy were supported in virtual voting centers around the city. The voting platform was extended to support cumulative voting by distributing 10 points over at least 3 projects [30, 43]. Voters are securely verified in a privacy-preserving way by providing a unique (hash) code that was generated by the city council using the social security number of each citizen (AHV card). The same platform was used for project campaigning by using the pairwise project comparisons feature. The winning projects were calculated using the method of equal shares and the Add1U completion method implemented within the Pabutools (the ones of Świecie and Wieliczka used the Add1 completion method, while Add1U would yield 1 and 3 more winners) [37]. Voting outcomes with the utilitarian completion method were also calculated and assessed in the exit survey (Table LABEL:table:exit-voting_outcome).
4.2 Evaluation of legitimacy
To assess the legitimacy of ballot aggregation methods, we rely on well-established theory of political legitimacy [34] and social choice [44, 45]. Among the three notions of input (voting participation), throughput (the voting process) and output (voting outcome) legitimacy [46, 47], we mainly focus in Section 2.2 on output legitimacy. However, we also assess the choice of ballot aggregation method before and after voting with equal shares as well as the effect of explaining equal shares, both capturing throughput legitimacy.
As such, by legitimacy we refer to voters who remain satisfied and accept the voting outcome of a ballot aggregation method (explained or not) even if they are among the losers, i.e. the did not vote for a winning project [48]. By considering all 24 combinations of winner vs. loser voters who are satisfied vs. unsatisfied (see Figure 3), we have a rigorous full factorial design to assess the legitimacy of ballot aggregation methods: which method voters prefer (Question MB.4 and MB.7 in Table LABEL:table:entry-voting_input), find fairer (Question MB.6 and MB.9) and more convenient (MB.10).
To distinguish winner and loser voters who are satisfied by the voting outcome from those who are not, we combine (i) satisfaction proxies based on survey questions with (ii) actual measurements of voters’ representation by the voting outcomes.
Proxies of satisfied vs. unsatisfied voters. We use different satisfaction proxies based on questions of the exit survey in Table LABEL:table:entry-digital (OB.1). In particular, we assess whether voters are satisfied with the voting outcome of City Idea (OB.1.1), whether they accept this outcome (OB.1.2) and whether they feel the outcome accurately represents the will of Aarau citizens (OB.1.7). Based on 11-point likert-scale responses, we split the populations two quartiles, the first one for the satisfied voters and the second one for the unsatisfied ones.
Determining winner vs. loser voters. To distinguish winner and loser voters by the voting outcome of equal shares, we measure the mean representation of voters’ preferences in the voting outcomes. The representation of a voter under cumulative voting is determined by the total points assigned by a voter to projects that win over the total points the voter assigns to projects. Winners represent the first quartile, while losers the second one. We also measure the representation of voters’ preferences by exclusive winning projects, which are the ones that win using the equal shares, but loose using the standard method. Exclusive representation reflects more prominently for voters the representation gain of equal shares over the standard method. We also distinguish voters to those who assigned points to the popular loser project of ‘Wild bees’ paradise’, which was the 3rd most popular project that is not elected under equal shares but it would be elected under the standard method. This allows us to assess the legitimacy of a particular effect of equal shares: ‘sacrificing’ a more expensive popular project, to elect instead several other projects that increase in overall the proportional representation of voters. Whether supporters of such exclusive losing project, which is popular and expensive, can accept this compromise for public good is a key question we explore in Section 2.2.
4.3 Explaining shifts between ballot aggregation methods
In this section, we present a causal analysis of how personal traits influence shifts in individual voter preferences for aggregation methods, based on the pre-voting survey, the actual voting process and the post-voting survey.
Prediction modeling for causal explanations. We explore a broad spectrum of factors (independent variables) that explain shifts in the choices of the ballot aggregation methods (dependent variable) as illustrated in Figure 3 and Section 4.2. The independent variables include personal human traits that can be found in the surveys related to political engagement and interest, altruistic behavior, voting outcome preferences, project preferences, and demographics. Different prediction models are constructed for shifts with and without method explanations. The shifts between methods model the dependent variable A and they are represented by a four-class classification problem, capturing transitions between aggregation methods: (i) standard to equal shares, (ii) standard to standard, (iii) equal shared to standard and (iv) equal shares to equal shares.
Prediction algorithms. Prediction models are constructed using tree-based gradient boosting algorithms [49] and neural networks with three hidden layers [50]. Different groups of personal human traits are assessed for the capacity to causally explain shifts between different ballot aggregation methods. The selection of these algorithms aligns well with how they handle categorical data (e.g. CatBoost [51] and neural networks [52] known to perform well even on smaller datasets). Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the models are illustrated in Section LABEL:sec:shifts-methods.
Approach for causal explanations. Once predictions are made, we derive the relative importance of the personal traits via a two-step explainable framework. Initially, an ablation study [53] is performed to systematically evaluate the relative predictive contribution of each personal human trait. In the second step, we employ model-agnostic Shapley Additive Explanations (SHAP) [54] to assess the individual contribution of each trait to the prediction performance. More details regarding the performance of the explanatory framework can be found in Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-lossLABEL:tab:trait-survey-map and Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
Data Availability
All collected and generated data along with the software code are available at:
http://doi.org/10.5281/zenodo.15162222 (with restricted access until publication).
Code Availability
The improved Stanford participatory budgeting platform is available at:
https://github.com/TDI-Lab/SPB-Stadtidee.
Author Contributions
EP: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Resources, Writing - Original Draft, Visualization, Supervision, Project administration, Funding acquisition. SM: Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Review & Editing, Visualization. TW: Software, Validation, Investigation, Data Curation, Writing - Review & Editing. JCY: Conceptualization, Methodology. FBH: Software, Validation, Data Curation. RHF: Conceptualization, Methodology, Data Collection, Investigation, Resources, Supervision, Writing - Review & Editing, Project administration, Funding acquisition DH: Conceptualization, Investigation, Resources, Writing - Input to Original Draft, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
Competing Interests
The authors declare no competing interests.
Additional Information
Supplementary information: The online version contains supplementary material available at https://doi.org/. Correspondence and requests for materials should be addressed to Evangelos Pournaras.
Acknowledgments
This project received support by the SNF NRP77 ‘Digital Transformation’ project "Digital Democracy: Innovations in Decision-making Processes", #407740_187249 and the UKRI Future Leaders Fellowship (MR/W009560/1): ‘Digitally Assisted Collective Governance of Smart City Commons–ARTIO’. The authors would like to thank the representatives from the city of Aarau for supporting and running the "Stadtidee" [translated: "City Idea"] campaign. The authors would also like to thank Abhinav Sharma for supporting the team at University of Leeds to collect the voting data as well as Michael Buess from Demoscope for his support with the surveys. The authors are grateful to Lodewijk Gelauff and Ashish Goel for their invaluable support they provided us to extend the Stanford Participatory Budgeting platform.

### v2

@@
 11footnotetext: Corresponding author: Evangelos Pournaras, School of Computing, University of Leeds, Leeds, UK, E-mail: e.pournaras@leeds.ac.uk
-Voting methods are instrumental design element of democracies.
+Voting methods are instrumental design elements of democracies.
 Citizens use them to express and aggregate their preferences to reach a collective decision.
 However, voting outcomes can be as sensitive to voting rules as they are to people’s voting choices.
-Despite the significance and inter-disciplinary scientific progress on voting methods, several democracies keep relying on outdated voting methods that do not fit modern, pluralistic societies well, while lacking social innovation.
+Despite significance and interdisciplinary scientific progress, several democracies keep relying on outdated voting methods that do not fit modern, pluralistic societies well, while lacking social innovation.
 Here, we demonstrate how one can upgrade real-world democracies, namely by using alternative preferential voting methods such as cumulative voting and the method of equal shares designed for a proportional representation of voters’ preferences.
-By rigorously assessing a new participatory budgeting approach applied in the city of Aarau, Switzerland, we unravel the striking voting outcomes of fair voting methods: more winning projects with the same budget and broader geographic and preference representation of citizens by the elected projects, in particular for voters who used to be under-represented, while promoting novel project ideas.
-We provide profound causal evidence showing that citizens prefer proportional voting methods, which possess strong legitimacy without the need of very technical specialized explanations.
+We rigorously evaluate the striking voting outcomes of these fair voting methods in a new participatory budgeting approach applied in the city of Aarau, Switzerland, including past and follow-up evidence.
+Results show more winning projects with the same budget.
+They also show broader geographic and preference representation of citizens by the elected projects, in particular for voters who used to be under-represented.
+We provide causal evidence showing that citizens prefer proportional voting methods, which possess strong legitimacy without the need of very specialized technical explanations.
 We also reveal strong underlying democratic values exhibited by citizens who support fair voting methods such as altruism and compromise.
-These findings come with a global momentum to unleash a new and long-awaited participation blueprint of how to upgrade democracies.
+These findings come with the momentum to unleash a new and long-awaited participation blueprint of how to upgrade democracies globally.
 1 Introduction
-Reaching collective decisions that benefit all often remains an unfulfilled promise.
-While voting methods have been instrumental to adaptively govern societies, they are often based on outdated approaches to foster collective intelligence [1, 2, 3].
-This may put democracies at risk instead of empowering them [4].
-Over-relying on limited-choice majoritarian voting often results in promoting competition instead of cooperation, in marginalization of minorities instead of inclusion, in few winners and many losers [5, 4, 6, 7, 2, 3, 8].
-Polarization, exacerbation of inequalities, and radicalization of the public sphere [9, 10] are prominent implications of voting methods that fail to foster fairness by design.
+Reaching collective decisions that benefit the broader society often remains an unfulfilled promise.
+While voting methods have been instrumental to govern societies with citizens’ participation, they are often based on outdated approaches to foster collective intelligence [Kearns2009, Arnold2017, Mann2017].
+This may put democracies at risk instead of empowering them [Emerson2020].
+Standard majoritarian voting methods, for example, over-prioritize popular preferences over those of under-represented groups.
+Disregarding alternative voices and their value often results in lost opportunities for all.
+Over-relying on limited-choice majoritarian voting frequently results in promoting competition instead of cooperation, in marginalization of minorities instead of inclusion, in few winners and many losers [Santos2024, Emerson2020, Pournaras2020, Helbing2023, Arnold2017, Mann2017, Hanggli2021].
+Polarization, exacerbation of inequalities, and radicalization of the public sphere [Lorenz2023, Bruter2017] are prominent implications of voting methods that fail to foster fairness by design.
+Therefore, the lack of fair, proportional representation of the different preferences in a society is a persistent setback to address.
 The challenge of adopting democratic innovations.
-Proportional voting methods and collective decision making have attracted broad interest by inter-disciplinary research communities [11, 12, 13].
-This includes the fields of social choice, computational social science, economics, complexity theory and political science.
-Nevertheless, adopting the developed methods in practice and changing the status quo is itself a long-standing and multi-faceted challenge [14].
-It often comes as a ‘chicken and egg’ problem: difficulties in communicating alternative voting methods effectively and lack of evidence of how they function in real world cause risk aversion at the political level and limited will for change, which further maintains the lack of evidence, also in regards to the legitimacy perception of voters [15, 16, 17, 2, 10].
-Such unprecedented complexity for democratic innovations requires rigorous scientific methods [17, 14, 10, 18].
+Proportional voting methods and collective decision making have attracted a broad interest by interdisciplinary research communities [Pukelsheim2017, Lewis2018, Fisher2018].
+This includes the fields of social choice, computational social science, economics, complexity theory, and political science.
+Nevertheless, adopting these methods in practice and changing the status quo is itself a long-standing and multi-faceted challenge [Madhavan2017].
+It often comes as a ‘chicken and egg’ problem, (i) with difficulties in communicating alternative, fairer voting methods effectively and (ii) with a lack of evidence of how they function in the real world, which mutually enforce each other.
+This causes risk aversion at the political level and limited will for change, which further maintains the lack of evidence, also with regard to the legitimacy perception of voters [Cho2018, Park2023, Noveck2017, Arnold2017, Bruter2017].
+Such unprecedented complexity for democratic innovations requires rigorous scientific investigation [Noveck2017, Madhavan2017, Bruter2017, Eliassi2020].
 A blueprint for shifting to fairer voting methods is urgently needed.
@@
 These include citizen assemblies, government labs (“GovLabs"), open innovation (open access/data/source, creative commons), crowd sourcing/sensing/funding, hackathons, fablabs, makerspaces, and citizen science.
-While related to some of these recent trends, our article focuses on significant interventions in participatory budgeting.
-This is a process designed to distribute a public budget of a city council in a bottom-up way by engaging citizens in collaborative brainstorming sessions to propose their own projects, solutions, and ideas, on which they later vote [19].
-Citizens may also be involved in the implementation of selected projects.
-Participatory budgeting offers a great potential for demonstrating grass-roots democratic innovations in real-world, which can be the blueprint to scale up further in other democratic processes.
-Started in Porto Alegre in 1989, participatory budgeting has since been adopted in many places across the globe, from remote communities in Kenya, Indonesia, and the Philippines to major cities such as Barcelona, Warsaw, Helsinki, Paris, and New York [20, 21].
-Its uptake has been particularly significant in Europe, with a surge in participatory initiatives in recent years [19].
-By 2019, Europe accounted for more than half of all participatory budgeting schemes worldwide, with more than 5,000 initiatives [21].
-Despite this uptake, the impact of fairer voting methods on voting outcomes as well as project ideation and implementation are, so far, under-explored questions.
-This is despite their profound importance for the quality of the voting outcomes.
+While related to some of these recent trends, our article focuses on novel interventions in participatory budgeting and their broader impact.
+This is a process designed to distribute some public budget of a city council in a bottom-up way by engaging citizens in collaborative brainstorming sessions to eventually propose their own projects, solutions, and ideas, on which they later vote [Fairstein2023].
+Citizens may also be involved in the implementation of the selected projects.
+How participatory budgeting generalizes.
+Participatory budgeting offers a great potential for demonstrating grassroots democratic innovations in the real world, which can be a blueprint that can be scaled up and wide later in other (collective) decision-making processes.
+For instance, the selection of projects to counter environmental problems, the distribution of research funds, the use of pensions funds, and entrepreneurial or stock market investments are all timely and pertinent budgeting processes that can benefit remarkably from budget-aware voting methods and, hence, from democratic innovations.
+Initiated in Porto Alegre back in 1989, participatory budgeting has since been adopted in many places across the globe, from remote communities in Kenya, Indonesia, and the Philippines to major cities such as Barcelona, Warsaw, Helsinki, Paris, and New York [Wampler2021, Dias2018].
+Its uptake has been particularly significant in Europe, with a surge of participatory initiatives in recent years [Fairstein2023].
+By 2019, Europe accounted for more than half of all participatory budgeting schemes worldwide, with more than 5,000 initiatives [Dias2018].
+Despite the profound importance of voting methods for the quality of the outcomes [Arrow1964], the impact of fairer voting methods has, so far, been under-explored.
 Collective choice made fairer.
-We demonstrate a breakthrough for democracies by applying two alternative voting methods designed for a more expressive and proportional representation of citizens in a real-world setting.
-This has been achieved with a novel field test designed for a systematic assessment of a new participatory budgeting process in the city of Aarau in Switzerland.
-We used the ballot format of cumulative voting [22, 23] allowing voters to distribute 10 points over at least 3 out of a total of 33 project ideas.
-Compared to other prominent input voting methods such as approval, ranking, or knapsack voting [24], cumulative voting enables voters to express their preferences more effectively by weighting their support to different options.
+We demonstrate a breakthrough for democracies by applying alternative voting methods designed for a more expressive and proportional representation of citizens in real-world settings.
+This has been achieved with a novel approach for field-testing and systematic assessing a new participatory budgeting process in the city of Aarau, Switzerland, called “Stadtidee” (“City Idea”).
+The participatory budgeting process there was so successful that it is planned to repeat.
+It was also followed by similar implementations elsewhere (e.g.
+in Poland), which support our innovative approach and conclusions.
+The outcomes in Aarau are particularly stunning as compared to a similar “Stadtidee” format in Zurich, where the conventional majoritarian voting method was applied (see Section 2.2).
+Therefore, our field test is documented as a methodological breakthrough that is of real-world political relevance.
+We used the ballot format of “Cumulative Voting” [Cooper2007, Skowron2025] allowing voters to distribute 10 points over at least 3 out of 33 project ideas.
+Compared to other prominent input voting methods such as approval, ranking, or knapsack voting [Goel2019], cumulative voting enables voters to express their preferences more effectively by weighting their support for different options.
 This expression of preferences is more inclusive and can provide a more proportional representation of minority groups.
-Complementary, we used the voting method of equal shares [25] to aggregate the points allocated by citizens and elect the winning project ideas.
-Equal shares promotes a more proportional representation in comparison with majoritarian voting.
-This is favorable, as majoritarian approaches to elect winners can sometimes lead to a "tyranny of the majority", where "the winner takes it all", i.e.
-monopolizes all the budget [26].
-Proportional representation is achieved by assigning to every voter a decision power, which is given by an equal share of the budget.
-This manifests both in meeting voters’ preferences and in the geographical coverage of the winners [2].
-In practice, the method of equal shares is resourceful, as it is likely to select a larger number of projects, which are cost-efficient, while satisfying the preferences of more people and increasing overall voter satisfaction [27, 28].
-Given the applied principle of proportionality, we refer to these alternative decision-making approaches here as ‘fair voting methods’.
-As it turns out, they also promote legitimacy and inclusion.
-Our study provides universal findings, beyond Aarau, as we also study a large number of past participatory budgeting processes and recent follow ups adopting fair voting methods.
+Complementary, we used the “Method of Equal Shares” [Peters2021] to aggregate the points allocated by citizens and elect the winning project ideas.
+This voting method promotes a more proportional representation in comparison with majoritarian voting.
+Proportionality is the favorable aspect of fairness we study in this article.
+It avoids an undesirable feature of majoritarian voting approaches, which can sometimes lead to a "tyranny of the majority", where "the winner takes it all", i.e.
+monopolizes all the budget [Nyirkos2018].
+In contrast, proportional representation is achieved by assigning to each voter the same decision power, which is given by an equal share of the budget to be distributed.
+This manifests both, in a better reflection of voter preferences and in a fairer geographical distribution of winning projects [Arnold2017].
+In practice, the method of equal shares is resourceful, as it is likely to select a larger number of projects, which are cost-efficient, while satisfying the preferences of more people and increasing overall voter satisfaction [Maharjan2024, Yang2024] as shown in Section 2.1.
+Given the principle of proportionality that we applied, we refer to these alternative decision-making approaches here as “fair voting methods”.
+As it turns out, they also promote legitimacy and democratic values (see Section 2.3 and 2.4 respectively).
+Our study provides universal findings, beyond Aarau, as we also study a large number of past participatory budgeting processes and recent follow-ups adopting fair voting methods.
 Overcoming barriers of shifting to fair voting methods.
-Combining cumulative voting and equal shares in a real-world participatory budgeting process allows us to unravel: (i) how to systematically adopt fair voting methods.
-Apparently, for policy-makers and a general public familiar with elections, the method of equal shares is pretty different from the widespread and longstanding standard method of utilitarian "greedy" voting, which simply sums up the votes and selects the most popular project ideas fitting the budget in a majoritarian way; (ii) how the method of equal shares influences participation in the project ideation phase, allowing for the consideration of additional criteria during the voting process.
-Project costs become instrumental, as more expensive projects require broader support.
-This may discourage expensive projects, which could be compensated by a larger number of more resource-efficient projects which are fairer distributed in space [27]; (iii) the legitimacy of the voting outcomes.
-Although the method of equal shares is designed to promote fairness, sacrificing popular expensive projects may be perceived as unfair by their supporters, who may be a considerable share of the population.
-Whether the selection of more alternative projects can compensate for the potential disappointment is key for the perceived legitimacy of fair voting methods; (iv) causal factors that promote a shift to fair voting methods.
-The crucial question is, whether there are any personal traits or societal values that support and accelerate democratic upgrades based on fair voting methods.
+Combining cumulative voting (for casting the vote) and equal shares (for aggregating the points and electing the winning projects) in a real-world participatory budgeting process allows us to unravel:
+  • (i)
+    how to systematically adopt fair voting methods; apparently, for policy-makers and a general public familiar with elections, the method of equal shares is pretty different from the widespread and longstanding standard method of utilitarian “greedy” (i.e., majoritarian) voting, which simply sums up the votes and selects the most popular project ideas fitting the budget in a majoritarian way;
+  • (ii)
+    how the method of equal shares influences participation in the project ideation phase, allowing for the consideration of additional criteria during the voting process; importantly, project costs become instrumental, as more expensive projects require broader support, which may discourage expensive projects; this allows for a larger number of more resource-efficient projects which are fairer distributed in space [Maharjan2024];
+  • (iii)
+    the legitimacy of the voting outcomes; although the method of equal shares is designed to promote fairness, sacrificing popular expensive projects may be perceived as unfair by their supporters, who may be a considerable share of the population; whether the selection of more alternative projects can compensate for the potential dissatisfaction is key for the perceived legitimacy of fair voting methods;
+  • (iv)
+    causal factors that promote a shift to fair voting methods; the crucial question is, whether there are any personal traits or societal values that support and accelerate democratic upgrades based on fair voting methods.
 A systematic and rigorous approach to upgrade democracies with fair voting methods.
-We co-designed the participatory budgeting process of "Stadtidee" ("City Idea") in collaboration with the city of Aarau in Switzerland.
-It was conducted in 2023 with an ongoing project implementation.
-It is characterized by the following innovations (see Section 4.1 for more information): (i) the real-world application of the method of equal shares in combination with cumulative voting as the ballot format.
-Taken together, they enable a better, proportional representation of citizens’ preferences; (ii) an inclusive participation fostered throughout the entire process.
-The citizens proposed 161 ideas, out of which 33 were voted on by 1703 citizens, women in majority and children included, with 17 projects selected for implementation [29], see Table 1 for an outline; (iii) a survey study was conducted before and after voting.
-This allows us to rigorously assess the new process in depth over its entire lifetime.
-We are also able to assess the legitimacy of fair voting methods before and after voting and to profoundly understand the dynamics of voting outcomes that these methods generate, including causal and explanatory factors; (iv) a secure and privacy-preserving voting using and extending the Stanford open-source platform for participatory budgeting [30, 31].
-This ensures transparency and trust when using fair voting methods.
-These four aforementioned innovations offer a strong foundation of validity to rigorously test fair voting methods.
-They provide a blueprint for testing upgrades of democracy, which have been missing so far.
+We co-designed the participatory budgeting process called “Stadtidee” (“City Idea”) in collaboration with the city of Aarau, Switzerland.
+The process was implemented in the year 2023.
+It was characterized by the following innovations (see Section 4.1 for more information):
+  • (i)
+    a real-world application of the method of equal shares in combination with cumulative voting as the ballot format; combined with each other, these methods enable a better, proportional representation of citizen preferences;
+  • (ii)
+    an inclusive participation fostered throughout the entire process; the citizens proposed 161 ideas, out of which 33 were voted on by 1703 citizens, women in majority and children included, with 17 projects selected for implementation [CityIdeaReport2025], see Table 1 for an outline;
+  • (iii)
+    a survey study was conducted before and after voting with a control condition; this allowed us to rigorously assess the new process in depth over its entire lifetime; we were also able to assess the legitimacy of fair voting methods before and after voting and to profoundly understand the dynamics of voting outcomes that these methods generate, including causal and explanatory factors;
+  • (iv)
+    a secure and privacy-preserving voting using and extending the Stanford open-source platform for participatory budgeting [SPBStadtidee2023, Wellings2023]; this ensures transparency and trust when using fair voting methods.
+A comprehensive spatio-temporal evidence. To confirm and generalize the qualities of voting outcomes reached by fair voting methods, we highlight the continuum of evidence we provide:
+  • (i)
+    Local comparison: The actual use of equal shares in Aarau vs. the hypothetical use of the standard majoritarian method in Aarau (shown in Section 2.1);
+  • (ii)
+    Global comparison: The actual use of equal shares in Aarau vs. the actual use of the standard majoritarian method in past elections (shown in Section 2.2);
+  • (iii)
+    Confirmation on earlier evidence: The hypothetical use of equal shares in past elections vs. the actual use of the standard majoritarian method in past elections (shown in Section 2.2);
+  • (iv)
+    Confirmation on follow-up evidence: The actual use of equal shares in Aarau vs. the actual use of equal shares in follow-up elections of Aarau (shown in Section 2.2);
+These four aforementioned innovations, in combination with comprehensive spatio-temporal comparisons with earlier and follow up real-world elections, enabled a rigorous test of fair voting methods and offers a strong foundation for the validity of our conclusions.
+This provides a promising blueprint for innovative upgrades of democracies in the future.
 2 Results
-This articles presents three key results:
-1. Fair ballot aggregation results in voting outcomes with more winners, while achieving better representation of voters’ preferences and novel project ideas, also in space.
-2. Voters find the adopted ballot aggregation method of equal shares more legitimate, by preferring it over the majoritarian standard method and finding it fairer.
-3.
-The shift to a fairer voting method distinguishes supporting voters with strong democratic values from voters without these values.
-Democratic values that explain the preference for more proportional voting methods include altruism, opposition to costly popular projects, and support for more expressive and inclusive ballot formats.
-2.1 Fair ballot aggregation has distinguishing multi-faceted qualities
+This article presents three key results:
+  • (1)
+    Fair ballot aggregation results in voting outcomes with more winners, while achieving better representation of voter preferences and diversity of project ideas, also in space (Section 2.1 and 2.2).
+  • (2)
+    Voters find the adopted ballot aggregation method of equal shares more legitimate, preferring it over the majoritarian standard method and finding it fairer (Section 2.3).
+  • (3)
+    The shift to a fairer voting method benefits from democratic values and promotes them.
+In fact, democratic values such as altruism, support for more expressive and inclusive ballot formats, and opposition to costly popular projects causally explain the preference for more proportional voting methods (Section 2.4).
+2.1 Fair voting methods show multi-faceted qualities
 Why the method of equal shares results in more winners.
-Figure 1a shows the 33 projects proposed for voting and the related voting results.
-The following key observations can be made: (i) Equal shares selects 17 winning projects, with more than half of the proposed projects winning.
-In contrast, the standard utilitarian greedy method would select 7 projects only, which would imply 26 losing projects, thereby creating a significant deficit in representing the broad range of citizen preferences.
-(ii) The project entitled ‘Wild bees’ paradise’ is an example of an expensive popular loser using the method of equal shares, which would win under the standard method.
-With 33.64% of voters selecting this project, it is the 3rd most popular project, which would consume 40% of the overall available budget.
-The method of equal shares selects several other, lower-cost projects instead, each of which may be less popular, but they address together the preferences of a broader population.
-(iii) The mean budget share of the proposed projects is 14.33%, which is 48.37% lower than a typical proposed project within 345 past participatory budgeting elections conducted with the standard method [32].
-This illustrates the impact of ballot aggregation design already in the project ideation phase: it is more resourceful by encouraging lower-cost projects of different and often novel nature, with the potential for synergetic impact [27].
-Figure 1: Compared to the standard utilitarian greedy method, which is majoritarian in nature, the number of winning projects resulting from the method of equal shares increases dramatically, with more than half of the proposed projects winning, while citizens’ proposed project ideas are better represented geographically.
-(A) Equal shares elects 17 winning projects compared to 7 with the majoritarian standard method.
-The project ‘Wild bees’ paradise’, requiring 40% of the total budget, is popular winner under the standard method but loser under equal shares.
-This allows more low-cost projects to win, which, in combination, better represent voters’ preferences.
-(B) Under equal shares, five districts receive more winning projects as compared to the standard method, thereby strengthening the periphery of the city.
-(C) Spatial fairness increases by 44% under equal shares as measured by the relative reduction in the index of dispersion for the number of winning projects relative to the proposed ones for each district.
-Equal shares improves voters’ geographic representation, especially for city peripheries.
-Figure 1b illustrates that the method of equal shares results in 5 districts (27.77%) with more winning projects compared to the majoritarian standard method, while no district receives a lower number of projects.
+Figure 1a shows the 33 projects proposed for voting (very left column) and the related voting results (right columns under "Standard" and "Equal Shares").
+The following key observations can be made:
+  • (i)
+    Equal shares selects 17 winning projects, with more than half of the proposed projects winning; in contrast, the standard majoritarian utilitarian greedy method would select 7 projects only, which would imply 26 losing projects, thereby creating a significant deficit in representing the broad range of citizen preferences;
+  • (ii)
+    The project entitled “Wild bees’ paradise” is an example of an expensive popular loser using the method of equal shares, which would win under the standard majoritarian method; with 33.64% of voters selecting this project, it is the 3rd most popular project, but it would consume 40% of the overall available budget.
+Similarly, the project“Open Children Studio" is the other expensive popular loser consuming 18.2% of the budget.
+The method of equal shares selects several other, lower-cost projects instead, each of which may be less popular, but together they better address the preferences of a broader population: 75% representation compared to 50% that “Wild bees’ paradise” and “Open Children Studio" achieve together.
+  • (iii)
+    The mean budget share of the proposed projects is 14.33%, which is 48.37% lower than a typical proposed project within 345 past participatory budgeting processes conducted with the standard majoritarian method [pabulib]; this illustrates the impact of ballot aggregation design already in the project ideation phase: it is more resourceful by encouraging lower-cost projects of different and often novel nature.
+This has the potential for synergetic impact creation due to the greater variety of cost-effective projects [Maharjan2024].
+Figure 1: Compared to the utilitarian greedy ("Standard") method, which is majoritarian in nature, the number of winning projects resulting from the method of equal shares (see column "Equal Shares") increases dramatically, with more than half of the proposed projects winning, while the citizens’ proposed project ideas are better represented geographically.
+(A) Equal shares (s)elects 17 winning projects as compared to 7 with the majoritarian standard method.
+The project ‘Wild bees’ paradise’, requiring 40% of the total budget, is popular winner under the standard method, but a loser under equal shares.
+This allows more low-cost projects to win, which, in combination, represent the voter preferences better.
+(B) Under equal shares, five districts receive more winning projects as compared to the standard majoritarian method, thereby strengthening the periphery of the city.
+(C) Spatial fairness increases by 44% under equal shares.
+This is measured by the relative reduction in the index of dispersion for the number of winning projects relative to the proposed ones for each district.
+Equal shares improves the voters’ geographic representation, especially for city peripheries.
+Figure 1b illustrates that the method of equal shares results in 5 districts (27.77%) with more winning projects compared to the standard majoritarian method, while no district receives a lower number of projects.
 This particularly benefits the periphery of the city that would otherwise be under-represented.
-Moreover, Figure 1c measures the relative number of winning/proposed projects in each district.
+Moreover, Figure 1c measures the relative number of winning as compared to the number of proposed projects in each district.
 Using the method of equal shares, the spatial fairness increases by 44% when measured by the reduction in the index of dispersion, which demonstrates that more proposed project win in each district.
-These results confirm that the periphery of Aarau is better represented, both in terms of the voter preferences and the number of proposed project ideas selected.
+These results confirm that the periphery of Aarau is better represented by the method of equal shares, both in terms of the voter preferences and the number of proposed project ideas selected.
+2.2 Generalizing fair voting methods as a distinguishing paradigm
 The method of equal shares represents voters better.
-Figure 2a shows the share of voters for different cumulative levels of the mean representation of their preferences.
-Representation is measured by the relative number of projects that are preferred by a voter and win.
-Comparing the standard method of 345 past participatory budgeting processes [33, 32] with the outcomes for the equal shares method in Aarau, we observe 3.7% more voters with higher representation levels, in particular for lower levels of cumulative representation.
-This means that equal shares better preserves the representation of the population even for under-represented voters, whose preferences do not align well with the voting outcomes.
-Specifically, for the participatory budgeting process in Aarau, the voters’ representation is 19% higher under equal shares compared to the standard method and 14.5% higher on average compared to 345 earlier participatory budgeting processes.
-Figure 2: Compared to earlier participatory budgeting experience based on the standard utilitarian greedy method, the method of equal shares increases the representation of voters’ preferences and project categories, in both project ideation and selection phase.
-(A) The representation of voter preferences increases by 14.5% under equal shares compared to earlier participatory budgeting processes based on the standard utilitarian greedy method.
-(B) The projects proposed for voting under equal shares promote project categories that have traditionally been under-represented, such as welfare and health, while projects that usually come with high costs, such as public transit, lose appeal significantly.
-(C) It is also these otherwise under-represented projects that become the winning ones when using equal shares.
+Figure 2a shows the share of voters for different minimum levels of the mean representation of their preferences (i.e.
+what share of voters have at least 0.1, 0.2,…,0.9 representation).
+Representation is measured by the relative number of projects preferred (i.e.
+selected) by a voter among those that win.
+The mean is calculated first for the voters of each election and then among elections (Figure LABEL:fig:pabulib-representation-voters is generated by averaging all voters among all elections).
+This figure reveals whether we have few or many voters that are very satisfied or unsatisfied (high/low representation).
+To generalize our findings we compare the following:
+  • (i)
+    The standard method of 38/345 past participatory budgeting processes [Faliszewski2023, pabulib] (baseline) that share similar characteristics with the process in Aarau (matched, see Section LABEL:sec:choice-ballot-aggregation);
+  • (i)
+    The method of Equal shares applied hypothetically to these 38 past elections;
+  • (i)
+    The method of Equal shares as applied in Aarau.
+Figure 2: Compared to earlier participatory budgeting experiences based on the standard utilitarian greedy method, the method of equal shares improves the representation of voter preferences and project categories, both in the project ideation and the project selection phase.
+(A) 27.8 % more voters on average get any minimum representation level using the method of equal shares as compared to earlier participatory budgeting processes based on the standard utilitarian greedy method.
+(B) The projects proposed for voting under equal shares promote project categories that have traditionally been under-represented, such as welfare and health, while projects that usually come with high costs, such as public transit, lose appeal.
+(C) When using the method of equal shares, some otherwise under-represented projects are winning.
 On average, project categories show a 3% higher representation.
-Equal shares encourages proposed projects that are novel and win.
-Figure 2b illustrates the share of the proposed project categories of participatory budgeting in Aarau compared to the projects observed in earlier participatory budgeting processes [33, 32].
-We find a significant change in the nature of the proposed projects: welfare and health projects, which were usually under-represented, are more prominently featured, while public transit projects are less prominent.
-This shift is determined by the costs of the projects belonging in the two categories: the average cost of public transit projects in the participatory budgeting process of Aarau amounts to 8% of the overall budget, while welfare-related projects amount to 12.25%.
-In earlier participatory budgeting processes, these numbers were 18.19% and 7.52%.
-Figure 2c presents the share of the winning project categories in the participatory budgeting process of Aarau for the method of equal shares and for the majoritarian standard method.
-It also shows the share in earlier participatory budgeting experience that relied on the standard method.
-Health and sport would be disadvantaged with the standard method as it would reduce their share in the winning set by 19.13% and 18.05%, respectively.
+Strikingly, much more voters (difference of 57.7% to 80.7%) are well-represented to a minimum extent with equal shares in Aarau, even when considering up to the minimum representation of 70% (Figure 2a).
+In contrast, a larger voters share for more than 90% representation is observed for the past elections using the standard majoritarian method.
+Such so high representation levels imply less capacity for compromise, leaving a larger part of the population with lower representation levels.
+This representation gain is not so high (difference of 49.6% to 56.7%) if the method of equal shares is applied to the past 39 past elections.
+This can be explained by (i) the different nature of the proposed projects when using equal shares, coming with lower cost (48.4% lower budget share) and (ii) the larger number of winning projects (23.8% relative increase in the number of winning projects).
+Table LABEL:tab:representation-gain also confirms that participatory budgeting processes with the project and budget characteristics (matching criteria) of Aarau result in the highest representation gain.
+This strengthens further the evidence for the approach of Aarau as a blueprint for democratic updates based on fair voting methods.
+The method of equal shares encourages proposed projects that are novel and win.
+Figure 2b illustrates the share of the proposed project categories in the participatory budgeting in Aarau compared to projects proposed in earlier participatory budgeting processes [Faliszewski2023, pabulib].
+We find a significant change in the nature of the proposed projects: welfare and health projects, which were usually under-represented before, are more prominently featured, while public transit projects are less prominent.
+This shift is reflected by the costs of projects in two categories: the average costs of public transit projects in the participatory budgeting process of Aarau amount to 8% of the overall budget, while welfare-related projects amount to 12.25%.
+In earlier participatory budgeting processes, these numbers were 18.19% and 7.52%, respectively.
+Figure 2c presents the share of the winning project categories in the participatory budgeting process of Aarau for the method of equal shares and for the standard majoritarian method.
+It also shows the share in earlier participatory budgeting experiences that relied on the standard method.
+Health and sports would be disadvantaged by the standard method, as it would reduce their share in the winning set by 19.13% and 18.05%, respectively.
+Regional comparison: Aarau [equal shares] vs.
+Zurich [standard].
+To further demonstrate the impact and validity of the innovative voting approach used in Aarau, we compare it with the previous participatory budgeting process in Zurich, which used the standard majoritarian voting method.
+Zurich (which is 38 km away from Aarau, i.e.
+located in rather close proximity) ran the “City Idea” format in 2021 [ZurichPBWeb, ZurichPBReport].
+Although data of individual voters are not available for analysis to us, a comparison based on aggregate measures is a particularly revealing ‘crash-test’ because of the following interesting facts:
+  • (i)
+    With a population size of 423,193>22,013, a larger pool of citizens can be attracted to participate in the “City Idea” process in Zurich as compared to Aarau;
+  • (ii)
+    The winning/proposed projects originated from all quarters of Zurich, which aimed to encourage a broad engagement by the citizens of Zurich in the participatory budgeting process: West: 16/33, North: 14/24, East: 14/29, South: 17/49;
+  • (iii)
+    The available budget for all regions was around 540,000 CHF, which was much more than the 55,000 CHF provided by the smaller city of Aarau.
+Despite those distinguishing factors, which are in favor of Zurich, the approach in Aarau turned out to be more effective than the one in Zurich.
+Even though Zurich had a larger size and a more diverse geographical origin of its proposed projects, the process attracted only the votes of 1,804 citizens, which was approximately the same number as in Aarau, corresponding to a 17 times higher turnout in Aarau.
+Moreover, many of the proposed projects in Zurich claimed the maximum amount of 10,000 CHF, while the average cost was 6.6% of the total budget, which is even lower than the 15% in Aarau.
+Accordingly, the projects in Aarau came with different cost modalities and higher diversity.
+Although the standard majoritarian voting method used in Zurich prioritized popularity, there were 7 projects in Zurich that were not selected because they each fell short by just 1 to 4 additional votes.
+This creates several origins of dissatisfaction for the voters in Zurich, in contrast to the process in Aarau, where the fair voting methods create several new ways for voters to have their preferred projects winning as showed in Section 2.1.
 Broader evidence beyond the case of Aarau.
-To strengthen further our findings, we also highlight the case of two other participatory budgeting elections that used equal shares during 2023 in two cities of Poland: (i) Świecie and (ii) Wieliczka (“"Green Million").
-In the first case that was a follow up of Aarau, 17/22 projects are selected for funding, while the majoritarian standard method would result in only 9 winning projects.
-The representation of voters’ preferences increased from 61% to 80%,.
-The top-2 project in voters’ approvals is not funded, which requires 36.6% of the budget and is approved by 24.4% of the population.
-In the second case, 30/64 projects are selected for funding, while the majoritarian standard method would result in 23 winning projects.
-The representation of voters preferences increased from 61% to 71%,.
-The top-2 project in voters’ approvals is not funded, which requires 10% of the budget and is approved by 8.9% of the population.
-Both cases confirm the distinguishing voting outcomes of equal shares: proportional representation is achieved by funding more projects that increase the overall satisfaction with compromises on popular expensive projects.
-In contrast to these two elections, further novel data are collected during the participatory budgeting process of Aarau that allow us to rigorously explore the legitimacy of fair voting methods and which citizens’ values explain the shift to those.
-2.2 Overcoming the adoption barrier of fair voting improves legitimacy
-Adopting a new ballot aggregation method: a rigorous evaluation.
+To strengthen our findings further, we also highlight the case of two other participatory budgeting processes that used the method of equal shares during 2023 in two cities of Poland: (i) Świecie and (ii) Wieliczka (“Green Million”).
+In the first case of Świecie that was a follow-up of Aarau, 17 of 22 projects are selected for funding, while the majoritarian standard method would have resulted in only 9 winning projects.
+In addition, the representation of voters’ preferences increased from 61% to 80% by the method of equal shares.
+The top-2 project in terms of voter approvals was not funded, which would have required 36.6% of the budget, while it was approved by 24.4% of the population.
+In the second case of Wieliczka, 30 of 64 projects were selected for funding, while the majoritarian standard method would have resulted in 23 winning projects.
+Here, the representation of voters’ preferences increased from 61% to 71%.
+The top-2 project in voters’ approvals was also not funded, which would have required 10% of the budget and was approved by 8.9% of the population.
+Both cases confirm the distinguishing voting outcomes we found for the method of equal shares: proportional representation is achieved by funding more projects, while a few expensive popular projects are sacrificed.
+However, this increases the overall satisfaction.
+In contrast to these two elections, further novel data were collected during the participatory budgeting process in Aarau, which allow us to rigorously explore the legitimacy of fair voting methods and the citizen values that explain the shift to those.
+2.3 Adopting fair voting improves legitimacy
+Rigorous evaluation of a new ballot aggregation method.
 Figure 3 illustrates which ballot aggregation method voters prefer and which they find fairer in the entry survey before voting and in the exit survey after voting.
 Section LABEL:sec:choice-ballot-aggregation also shows additional results of which method voters find more convenient.
-Section LABEL:sec:equal-shares-explanation further illustrates how the two methods are presented to voters, in particular the voting outcomes and the explanations, based on which the measurements are made.
-The shift from the majoritarian standard method to the method of equal shares becomes obvious by the gain in the share of voters, who prefer and find fairer equal shares after voting (rigorously measured on the same matched population).
-We also assessed which method voters found more convenient in Figure LABEL:fig:convenience.
-The legitimacy of the two ballot aggregation methods is determined by measuring the satisfaction of voters with the voting outcomes of each method (even if the voter classifies as a ‘loser’, based on different measurements), see Section 4.2 for more information.
-To assess legitimacy thoroughly, the voters’ gain is disaggregated further down to all 24 combinations of winners/losers and satisfied/unsatisfied voters (see Figure LABEL:fig:legitimacy_fairness_preference_heat).
-We group those into three groups in Figure 3 for simplicity: satisfied winners, unsatisfied losers and mixed, each consisting of 8 combinations.
+Section LABEL:sec:equal-shares-explanation further illustrates how the two methods, based on which our measurements were made, were presented to the voters, in particular the voting outcomes and the explanations.
+The shift from the majoritarian standard method to the method of equal shares becomes obvious by the gain in the share of voters, who prefer equal shares after voting and find it fairer (rigorously measured on the same matched population).
+We also assessed which method voters found more convenient (see Figure LABEL:fig:convenience).
+The legitimacy of the two ballot aggregation methods was determined by measuring the satisfaction of voters with the voting outcomes of each method.
+Remarkably, it increased even if the voter is classified as a ‘loser’ (based on different measurements); see Section 4.2 for more information.
+To assess legitimacy thoroughly, the voters’ gain (from before to after voting) was disaggregated further down to all 24 combinations of winners/losers and satisfied/unsatisfied voters (see Figure LABEL:fig:legitimacy_fairness_preference_heat).
+For simplicity, in Figure 3 we grouped those into three groups: satisfied winners, unsatisfied losers and mixed, each consisting of 8 combinations.
 Three ways to measure each of the winners/losers and satisfied/unsatisfied are illustrated in Section 4.2.
-Figure 3: In the exit survey after voting, the method of equal shares is found more preferred and fairer than the majoritarian standard ballot aggregation method.
-Strikingly, before voting and adopting equal shares, there is no clear preference for any of the two methods, while the standard method is found fairer.
-Voting outcomes themselves are sufficient for individuals to determine equal shares as their preferred and fair ballot aggregation method.
-The explanation of the methods does not significantly contribute to shifting the choice to equal shares.
-The legitimacy of equal shares is assessed based on voters’ (A) preference and perceived (B) fairness of equal shares over the standard method.
-The x-axis determines the difference of the voters share between those choosing equal shares over the standard method.
-Legitimacy is assessed based on 24 combinations of winners/losers and satisfied/unsatisfied voters that are grouped into (i) winners and satisfied, (ii) losers and unsatisfied and (iii) mixed.
+Figure 3: In the exit survey after voting, the method of equal shares was found to be more preferred and fairer than the standard majoritarian ballot aggregation method.
+Strikingly, before voting and adopting equal shares, there was no clear preference for any of the two methods, while the standard method was found fairer.
+The voting outcomes themselves were sufficient for individuals to determine equal shares as their preferred and fairer ballot aggregation method.
+(The explanation of the methods does not significantly contribute to shifting the choice to equal shares.) Furthermore, the legitimacy of equal shares was assessed based on (A) the voters’ preference and (B) their perceived fairness of equal shares over the standard method.
+The x-axis determines the difference (gain) of the voters proportion between those choosing equal shares and those choosing the standard method.
+Legitimacy is assessed based on 24 combinations of winners/losers and satisfied/unsatisfied voters that are grouped into (i) satisfied winners, (ii) unsatisfied losers and (iii) mixed.
 Winners/losers are determined in three different ways: 1st and 2nd quartile of representation score for the winning projects that are (i) non-exclusive and (ii) exclusive between the two ballot aggregation methods.
-(iii) The non-supporters and supporters of the top popular losing project.
-Satisfied/unsatisfied voters are determined by three post-voting survey questions that assess: (i) satisfaction, (ii) acceptance of outcome and (iii) accurate representation of will.
-The adoption of equal shares exhibits a shift to be the most preferred and fairer method.
-The results of Figure 3 show that, in the exit survey after voting as compared to the entry survey before, there is a significant gain in voters who prefer the method of equal share and find it fairer.
-This is particularly prominent for unsatisfied losers who choose the standard method as fairer before voting but switch to equal shares after voting (see Figure LABEL:fig:legitimacy_fairness_preference_heat).
-More specifically, 54% of the population prefers equal shares before voting, while after voting, supporters of equal shares raise to 80% (p=0.0034).
-Similarly, 45% of voters find equal shares fairer before voting and 89% fairer after voting (p=0.028).
+(iii) Non-supporters and supporters of the top popular losing project.
+Satisfied/unsatisfied voters are determined by three post-voting survey questions that assess: (i) satisfaction, (ii) acceptance of the outcome and (iii) accurate representation of will.
+The adoption of the method of equal shares exhibits a shift to be the most preferred and fairer method.
+The results of Figure 3 show that, there is a significant gain in voters who prefer the method of equal share and find it fairer in the exit survey after voting when compared to the entry survey before.
+This is particularly prominent for unsatisfied losers who chose the standard method as fairer before voting but switched to equal shares after voting (see Figure LABEL:fig:legitimacy_fairness_preference_heat).
+More specifically, 54% of the population preferred equal shares before voting, while after voting, supporters of equal shares raised to 80% (p=0.0034).
+Similarly, 45% of voters found equal shares fairer before voting and 89% fairer after voting (p=0.028).
 The method of equal shares shows stronger legitimacy.
 The significant legitimacy of equal shares is reflected in several ways (see also Figure LABEL:fig:legitimacy_fairness_preference_heat): (i) Losers, determined as supporters of the top popular losing project (‘Wild bees’ paradise’), prefer the method of equal shares and find it fairer.
-(ii) Remarkably, the method of equal shares is preferred and found fairer even more among losers than among winning voters, who prefer the majoritarian standard method.
-(iii) The shift to equal shares after voting is even found among those who do not accept the voting outcome, which is a strong proxy of legitimacy [34, 35].
-When shift to equal shares is not achieved.
-However, we also found some cases where more voters shifted to the majoritarian standard method after voting, namely (i) satisfied voters preferring the standard method over unsatisfied voters preferring equal shares; (ii) voters accepting the voting outcome and finding the standard method fairer over voters who do not accept the voting outcome and find equal shares fairer; (iii) unsatisfied losers (or winners), who prefer the standard method over unsatisfied winners (or losers), who choose equal shares, when winners/losers are determined by the representation and their choice for the top popular losing project (‘Wild bees’ paradise’).
-See Figure LABEL:fig:legitimacy_fairness_preference_heat for a more disaggregated view.
+(ii) Remarkably, the method of equal shares is even found fairer and preferred to a higher extent among losing voters than among winning ones, who often prefer the standard majoritarian voting method.
+(iii) The shift to equal shares after voting is found fairer among those who do not accept the voting outcome, which is a strong proxy of legitimacy [Weatherford1992, Johnson2006].
+However, we also found some cases where more voters shifted to the standard majoritarian voting method after voting (see Section LABEL:sec:choice-ballot-aggregation).
 Voting outcome itself vs.
-method explanation: what matters.
-The method explanation has an additional contribution of 5.3% to the shift to equal shares as preferred method, while it has an addition contribution of 5.2% to the shift to equal share as fairer method.
-The method explanation has a small but statistically significant contribution to the shift to equal shares (with only exception the mixed setting for the preferred method, p=0.07).
-These results demonstrate that the voting outcome itself is sufficient to establish a change.
-2.3 Democratic values explain voters’ shift to fairer voting methods
-Towards fairer voting: A clash between democratic values and democratic backsliding.
-Figure 4 identifies distinguishing human factors that causally explain the shift from one ballot aggregation method before voting to another one after voting (with the explanation of method), as illustrated in Figure 3 and in Section 4.3.
+method explanation.
+The explanation of the voting method has an additional contribution of 5.3% to the shift towards equal shares as the preferred method, while it has an additional contribution of 5.2% to the shift towards equal share as the fairer method.
+Accordingly, the method explanation has a small but statistically significant contribution to the shift to equal shares (with the exception of the mixed setting for the preferred method, p=0.07).
+These results demonstrate that the voting outcome itself is sufficient to establish a change in favor of the method of equal shares.
+2.4 Democratic values explain voters’ shift to the fairer voting method
+Clash between democratic values and democratic backsliding.
+Figure 4 identifies human factors that causally explain the shift from one ballot aggregation method before voting to another one after voting (with the explanation of method), see also Figure 3 and Section 4.3.
 The classification model is illustrated in Section LABEL:sec:shifts-methods and comes with an accuracy of 85.92%.
-Strikingly, the group of voters who shift from the majoritarian standard method to equal shares exhibits democratic values that explain this transition, see Figure 4c.
-These voters come with distinguishing qualities (Figure 4b) such as being altruistic, compromisers and supporting more expressive, inclusive ballots.
+Strikingly, the group of voters who shift from the standard majoritarian voting method to equal shares exhibits democratic values that explain this transition, see Figure 4c.
+These voters come with distinguishing qualities (Figure 4b) such as being altruistic, being compromisers, and supporting more expressive, inclusive ballots.
 They are aware of the proposed projects and flexible to select low-cost projects, which also makes them more likely to oppose a popular costly project.
 This group is well-represented in the voting outcome.
-In contrast, the voters that shift from equal shares to the majoritarian standard method do not exhibit these democratic values, but are driven by self-interest, mistrust and dissatisfaction.
+In contrast, the voters that shift from equal shares to the standard majoritarian method do not show these democratic values.
+Instead they are driven by self-interest, mistrust and dissatisfaction.
 They support simpler ballot formats, are not well-represented in the voting outcomes, and support the popular loser project under equal shares.
-Figure 4: The shift to equal shares as the preferred method after voting is explained by distinguishing human traits that reflect the promotion of strong democratic values such as being altruistic, compromisers and supporting more expressive, inclusive ballots.
-In contrast, shifting to the majoritarian standard method is found among voters showing mistrust, under-representation, dissatisfaction, and self-interest.
-These voters also support the popular costly loser project of the equal shares and simple ballot methods.
+Figure 4: The shift to equal shares as the preferred method after voting is explained by distinguishing human traits that reflect the promotion of strong democratic values such as being altruistic, compromisers, and supporting more expressive, inclusive ballots.
+In contrast, shifting to the standard majoritarian method is found among voters showing mistrust, under-representation, dissatisfaction, and self-interest.
+These voters also support the popular costly loser project.
 A classification model (with 85.92% accuracy) is designed with four classes representing the preferred ballot aggregation method before and after voting (with explanation of methods).
@@
 These are prominently associated with democratic backsliding on the one hand, and promotion of democratic values on the other hand.
-(B) Summary of distinguishing qualities for those who prefer the method of equal shares over the majoritarian standard method after voting (feature difference of the two classes).
+(B) Summary of distinguishing qualities for those who prefer the method of equal shares over the standard majoritarian method after voting.
 Altruism, support of expressive, inclusive ballots and opposition to the popular, costly loser project of equal shares are qualities with the highest prominence for those voters preferring equal shares.
-(C) Human traits (features) that explain the preference of each voter before and after voting, in particular the transitions from the majoritarian standard method to equal shares and from equal shares to the standard method (see Figure 3).
+(C) Human traits (features) that explain the preference of each voter before and after voting, in particular the transitions from the standard majoritarian method to equal shares and from equal shares to the standard method (see Figure 3).
 The ‘*’ indicates that a feature is statistically significant (p<0.05) in explaining the voters’ choice of the method before and after voting.
 Distinguishing qualities of voters who shift preference to equal share.
-Figure 4b summarizes the distinguishing qualities of voters, who shift to equal shares by taking the difference of the human traits for the two transitions (standard to equal shares - equal shares to standard).
+Figure 4b summarizes the distinguishing qualities of voters, who shift to equal shares by taking the difference of the human traits for the two transitions (standard to equal shares – equal shares to standard).
 Altruism, opposition to the popular (costly) loser project and support of the expressive, inclusive ballot formats are the most prominent distinguishing qualities observed for voters who shift to equal shares.
@@
 A blueprint to upgrade democracies.
-In our study, we have shown how voting can better foster collective intelligence that is based on higher inclusion and diversity, with strong evidence provided by a rigorous participatory budgeting process in the city of Aarau in Switzerland.
-Subsequent participatory budgeting processes run using this approach further confirm our findings [36].
-In particular, we found that the majoritarian standard approach of voting had a number of deficits, such as the selection of a few projects, leaving a large portion of the project ideas unfulfilled.
-As the selected projects are implemented mainly in the city center, it largely fails to reach spatial fairness or inclusion.
-However, combining cumulative voting with the method of equal shares proves to be resourceful, as it promotes more novel winning projects, fairness, satisfaction, inclusion, trust and democratic values.
+In our study, we have shown how fair voting methods can better foster collective intelligence that is based on higher inclusion and diversity, with strong evidence provided by (i) a rigorous participatory budgeting process in the city of Aarau in Switzerland as well as (ii) extensive comparisons with other participatory processes.
+Subsequent processes using fair voting methods confirm our findings [equalsharesmethod].
+Accordingly, the standard majoritarian approach of voting has a number of deficits, such as the selection of few projects only, which often lack cost-effectiveness, leaving a large portion of good project ideas unrealized.
+As projects selected by majoritarian voting are typically implemented in the city center, they largely fail to achieve spatial fairness or inclusion.
+In the past, this issue has been addressed by dividing a city into multiple districts and running multiple budgeting processes in parallel, as in the case of Zurich in 2021.
+In contrast, combining cumulative voting with the method of equal shares proves to be resourceful, as it supports more winning projects.
+It also promotes fairness, satisfaction, inclusion, trust, and democratic values.
 Accordingly, this approach offers a powerful way of upgrading democracies.
-Implementing this approach, using digital means, promises to be a powerful, value-sensitive approach to promote digital democracy by design [7].
+Implementing this approach, using digital means, is also a value-sensitive approach to promoting digital democracy by design [Helbing2023].
 Success of project implementation.
-By the end of 2024, 9 of 17 projects selected by the "City Idea" process in Aarau were already implemented and 5 projects got an extension, while 3 projects could not be implemented mainly due to feasibility and budgetary constraints [29].
+By the end of 2024, 9 of 17 projects selected by the “City Idea” process in Aarau were already implemented and 5 projects got an extension, while 3 projects could not be implemented mainly due to feasibility and budgetary constraints [CityIdeaReport2025].
 These were among the top-11 projects.
-Two thirds of these projects would also have been winners if the majoritarian standard method had been used.
-These results demonstrate that having more winning projects using the equal shares method does not negatively affect the overall capacity of citizens to implement them.
+Two thirds of these projects would also have been winners if the standard majoritarian method had been used.
+These results demonstrate that having more winning projects using the method of equal shares does not negatively affect the overall capacity of citizens to implement them.
+In contrast, the standard majoritarian voting method applied in Zurich for the “City Idea” process did not produce so satisfactory results, and the press critically discussed the problem of cost-inefficient projects [ZurichPBReport].
 Participation levels.
 The participatory budgeting process in Aarau achieved significant participation of citizens at different levels.
-During the project ideation phase, 161 ideas were proposed, which exceeded the initial goal of the campaign to attract 100 ideas by far [29].
-The total of 1703 voters who participated is a significant number for such a campaign in a town with 14,310 residents having voting rights in 2023.
-Nevertheless, the turnout did not reach the level of official elections such as referenda and national elections.
-A 22% of the voters showed interest in participating, although they were not officially residents of Aarau, while 9% of the voters were not Swiss citizens.
+During the project ideation phase, 161 ideas were proposed, which exceeded by far the initial goal of the campaign to attract 100 ideas [CityIdeaReport2025].
+In comparison, 176 ideas were proposed and 135 of them presented for voting in Zurich in 2021 with a 19 times larger population and a 10 times larger total budget.
+The total of 1,703 participating voters in Aarau (as compared to 1,804 in Zurich) is a significant number for such a campaign in a town of 14,310 residents with voting rights in 2023.
+Nevertheless, the turnout did not reach the level of official elections such as referenda and national elections, yet.
+A 22% of voters showed interest in participating in the “City Idea” process even though they were not official residents of Aarau, while 9% of voters were not Swiss citizens.
 The average age of voters was 43, and there were more female than male voters.
-Children over 15 years old were allowed to vote.
-Among the low number of voters who abstained but participated in the surveys (296), the top-5 reasons for not voting were "no time" (17.1%), "forgetting" (12.8%), "lack of information" (9.9%), "unappealing projects" (7.7%) or some "other reason" (7.0%), see Question EB.7 in Table LABEL:table:exit-digital.
+Children over 15 years were allowed to vote.
+Among the low number of citizens who abstained but participated in the surveys (296), the top-5 reasons for not voting were "no time" (17.1%), "forgetting" (12.8%), "lack of information" (9.9%), "unappealing projects" (7.7%) or some "other reason" (7.0%), see Question EB.7 in Table LABEL:table:exit-digital.
 Upgrading democracies requires upgrading digital tools.
 We created and used a novel combination of open-source and value-sensitive digital tools to meet a broad range of expectations regarding transparency, security, privacy, reproducibility, explainability and fairness.
-This required a significant implementation effort, a mobilization of various communities, a strong commitment and the building of shared values with the city authorities, which is a process that has not been systematic pursued to this extent so far.
-A highlight of this effort was the privacy-preserving verification of citizens’ eligibility when voting with the social security number, which ensured a trusted voting procedure.
-The integration of our open-source implementation of cumulative voting [30] into the Stanford Participatory Budgeting platform is a stepping stone towards a more ambitious roadmap, which aims to make participatory tools more standardized and interoperable, in particular the ones supported by large open-source communities such as Decidim and Consul as well as the ones created by the computational social choice community such as Pabutools [37] and Preflib [38].
+This required a significant implementation effort, a mobilization of various communities, a strong commitment and the building of shared values with the city authorities, which is a process that has not been systematically pursued to this extent before.
+A highlight of this effort was the privacy-preserving verification of citizens’ eligibility when voting with the social security number, which ensured a trustworthy voting procedure.
+The integration of our open-source implementation of cumulative voting [SPBStadtidee2023] into the Stanford Participatory Budgeting platform is a stepping stone towards participatory tools that are more standardized and interoperable.
+This is also relevant for other large open-source communities such as Decidim and Consul as well as the ones created by the computational social choice community such as Pabutools [pabutools] and Preflib [Mattei2013].
 Explaining and adopting fair voting methods.
-A 53.17% of voters understood how equal shares calculates the winning projects (MB.3.3), while a 55.95% understood why popular projects are not easily funded by equal shares (MB.3.6), and 41.26% find that fair (MB.3.7).
+53.17% of voters understood how equal shares calculates the winning projects (MB.3.3), while 55.95% understood why popular projects are not easily funded by equal shares (MB.3.6), and 41.26% found this fair (MB.3.7).
 Strikingly, the results of Figure 3 demonstrate that explaining the expected difference in the voting outcomes reached by equal shares is sufficient for voters to prefer this method.
@@
 This result has significant implications for overcoming the long-standing hesitation with regard to democratic innovations, since citizens show a readiness to accept more complex decision-making methods as long as they understand the anticipated impact on the decision outcome.
-In addition, it helps when people experience the different result, which gives them a practical understanding of the method.
-Moreover, given that equal shares is designed for a proportional representation of voter preferences and is known for its stability [19], it is more unlikely to yield controversies over voting outcomes.
+In addition, the experience of the different results helps citizens to shape a practical understanding of the method.
+Moreover, given that equal shares is designed for a proportional representation of voter preferences and is known for its stability [Fairstein2023], it is more unlikely to cause controversies over voting outcomes.
 Nevertheless, the additional effort required to explain the method of equal shares should not be under-estimated.
 As shown in Section LABEL:sec:equal-shares-explanation, the introduction of equal shares requires a systematic communication effort over the lifetime of the participatory budgeting process, using creative media and visualizations suited for the broader public.
-At the project ideation phase, awareness of the method is required to mobilize different parts of the populations for ideas that may not have been considered under the majoritarian rule.
-As the costs of the proposed projects play a key role for selecting them under proportional representation, citizens need to be resourceful and prioritize different ideas, meaning a focus on cost-effective projects that get the support by a sufficiently large group of the population.
-Large and costly projects may be more risky to propose as they require the support of a broader population to get selected.
+At the project ideation phase, awareness of the method is required to mobilize the population for ideas that may not have been considered under the majoritarian rule.
+As the costs of the proposed projects play a key role for selecting them, citizens need to be resourceful and prioritize different kinds of ideas, which requires a focus on cost-effective projects that are supported by a sufficiently large group of the population.
+Instead, large and costly projects may be more risky to propose as they require the support of a broader population in order to get selected.
 Fair voting methods to foster a novel impact creation process.
 The influence of the chosen voting method on the proposed and selected project ideas can transform the process of creating real-world impact.
-For instance, novel welfare projects gain momentum, as they appeal to several interest groups, while they may also be easier to implement than costly infrastructural projects [27].
-We foresee that, on the long term, proportional voting methods could enable an incremental grass-roots implementation of sustainable development projects over several participatory budgeting campaigns, which could also support a fairer transition to a net zero future.
-Fair voting methods: a better-than-optimal paradigm.
+For instance, novel welfare projects gain momentum, as they appeal to several interest groups, while they may also be easier to implement than costly infrastructural projects [Maharjan2024].
+We foresee that, on the long term, proportional voting methods could enable an incremental grass-roots implementation of sustainable development projects over several participatory budgeting campaigns, which could also support a fairer transition to more sustainable future.
+Fair voting methods are a better-than-optimal paradigm.
 Over-simplifying, averaging or overriding people’s preferences while optimizing societies for goals that people may not even explicitly agree on is a significant risk that democracies face in the era of artificial intelligence (AI) run by Big Tech.
-In this context, the notion of ‘optimal’ solutions is actually misleading and displaced from the reality of people, as demonstrated by the inferiority of popular expensive projects or projects focused on city centers that are typically proposed in participatory budgeting processes under majoritarian standard voting.
+In this context, the notion of ‘optimal’ solutions is actually misleading and displaced from the reality of people, as demonstrated by the inferiority of popular expensive projects or projects focused on city centers that are typically proposed in participatory budgeting processes under standard majoritarian voting.
 In contrast, proportional voting methods work well for pluralistic settings, where voters may not agree on one and the same goal, but decisions are based on collective intelligence.
@@
 Small compromises, such as excluding a popular expensive project, open up opportunities for large collective returns with benefits for a great majority of people.
-For instance, the inclusion of under-represented health projects or of projects in the city periphery can reduce social tensions and benefit a broader population.
+For instance, the inclusion of otherwise under-represented health projects or projects in the city periphery can reduce social tensions and benefit a broader population.
 The demonstrated approach of combining cumulative voting with the method of equal shares delivers a powerful blueprint for upgrading democracies.
-It allows societies to achieve (i) a larger number of realized projects within the respectively available budget; (ii) a greater resourcefulness of novel projects and a multi-faceted impact; (iii) more participation and inclusion; (iv) greater transparency and trust in the voting process; (v) better representation of citizen preferences and higher satisfaction; (vi) more (spatial) fairness; (vii) higher legitimacy thanks to a greater acceptance of the collective outcomes; (viii) strengthened democratic values of voters.
-We refer these impacts "better than optimal" as they tend to be a win-win for all, while they are distinguished from what existing optimization and data-driven decision-making approaches typically offer.
+It allows societies to achieve:
+  • (1)
+    a larger number of realized projects within the available budget;
+  • (2)
+    a greater resourcefulness of novel projects and a multi-faceted impact;
+  • (3)
+    more participation and inclusion;
+  • (4)
+    greater transparency and trust in the voting process;
+  • (5)
+    better representation of citizen preferences and higher satisfaction;
+  • (6)
+    more (spatial) fairness;
+  • (7)
+    higher legitimacy thanks to a greater acceptance of the collective outcomes;
+  • (8)
+    strengthened democratic values of voters.
+We call these impacts “better than optimal” as they tend to be a win-win for all, while they are distinguished from what existing optimization and data-driven decision-making approaches typically offer.
 Fair voting methods in the era of artificial intelligence.
-We would like to end on a remark addressing the use of data-driven AI systems for decision making.
+We would like to complete this discussion on a remark addressing the use of data-driven AI systems for decision making.
 While we consider AI a very powerful technology with numerous promising application areas, how to use it in the context of governing democracies must be studied very carefully.
 AI safeguards for democracy are critical and yet to be created.
-Existing decision-making methods supported by AI solutions show alarming vulnerabilities and are still perceived as lacking fairness and legitimacy [39, 40, 41].
-For instance, biases and inconsistencies by unconsolidated use of large language models can exacerbate by voting methods with a deficit on fairness [42].
-It is not clear when and to what extent AI risks can be addressed.
+Existing decision-making methods supported by AI solutions show alarming vulnerabilities and are still lacking fairness and legitimacy [Saetra2022, Helbing2015, Koster2022].
+For instance, voting methods with a deficit on fairness can exacerbate biases and inconsistencies by unconsolidated use of large language models [Majumdar2024].
+So far, it is not clear how to address some existing AI risks.
 Therefore, alternative approaches such as the one presented here are urgently needed.
 4 Methods
-We illustrate the designed participatory budgeting campaign of City Idea, the evaluation approach of legitimacy and the models for causal explanations for the shifts between ballot aggregation methods before and after voting.
-4.1 The participatory budgeting campaign of City Idea: field test design
-Table 1 outlines the participatory budgeting campaign of City Idea.
+We illustrate the designed participatory budgeting campaign of “City Idea”, the evaluation approach of legitimacy and the models for causal explanations for the shifts between ballot aggregation methods before and after voting.
+4.1 The participatory budgeting process of “City Idea”: Field test design
+Table 1 outlines the participatory budgeting campaign of “City Idea”.
 It consists of six phases over the passage of nine months.
 An entry and exit survey were conducted before and after voting respectively.
-The entry survey was launched via physical invitation letters to all citizens sent by city council (10,000 A5 flyers [29]).
+The entry survey was launched via physical invitation letters to all citizens sent by city council (10,000 A5 flyers [CityIdeaReport2025]).
 This resulted in 3592 respondents.
 Some of these filled in their email to voluntarily participate in the exit survey and this resulted in 808 respondents.
-The surveys included questions about demographics and political opinion (Table LABEL:table:entry-demographics-LABEL:table:exit-demographics), engagement and digital skills (Table LABEL:table:entry-digital-LABEL:table:exit-digital), the City Idea voting (Table LABEL:table:entry-city_idea-LABEL:table:exit-city_idea), the voting outcome preference (Table LABEL:table:entry-voting_outcome-LABEL:table:exit-voting_outcome) and the voting input and aggregation method (Table LABEL:table:entry-voting_input-LABEL:table:exit-voting_input).
+The surveys included questions about demographics and political opinion (Table LABEL:table:entry-demographics-LABEL:table:exit-demographics), engagement and digital skills (Table LABEL:table:entry-digital-LABEL:table:exit-digital), the “City Idea” voting (Table LABEL:table:entry-city_idea-LABEL:table:exit-city_idea), the voting outcome preference (Table LABEL:table:entry-voting_outcome-LABEL:table:exit-voting_outcome) and the voting input and aggregation method (Table LABEL:table:entry-voting_input-LABEL:table:exit-voting_input).
 Survey questions encode variables, which are observed for changes before and after voting.
 The overall study has received ethical approval from University of Fribourg (#2021-680).
-Table 1: The rigorously designed participatory budgeting campaign of City Idea with 6 phases over a period of 9 months in 2023.
+Table 1: The rigorously designed participatory budgeting campaign of “City Idea” with 6 phases over a period of 9 months in 2023.
 Participants of the entry survey, voting and exit survey are linked using an anonymous identifier (hash) of the citizens’ social security number.
@@
 Project ideation, deliberation and campaigning.
-Citizens proposed more than 161 project ideas out of which 33 projects are selected to put for voting [29].
+Citizens proposed more than 161 project ideas out of which 33 projects are selected to put for voting [CityIdeaReport2025].
 Projects are eliminated or similar project ideas are combined to strengthen their chances to win.
@@
 It also organized two press releases that resulted in 17 media reports.
-Moreover, the interactive information campaign ‘know your projects’ was launched in the city that engaged 73 participants in 1682 random pairwise comparisons of projects (see Table LABEL:tab:pairwise_winners for further information) via the Stanford Participatory Budgeting platform  [43].
+Moreover, the interactive information campaign ‘know your projects’ was launched in the city that engaged 73 participants in 1682 random pairwise comparisons of projects (see Table LABEL:tab:pairwise_winners for further information) via the Stanford Participatory Budgeting platform  [Wellings2023].
 Awareness about the method of equal shares and what projects are expected to win was raised, emphasizing the challenge of electing projects that cover a large share of the budget.
@@
 The city of Aarau has a total of 22,593 citizens, out of which 17,428 are Swiss citizens.
-A total of 1703 voters participated in the City Idea campaign, out of which 1330 are citizens of Aarau.
+A total of 1703 voters participated in the "City Idea" campaign, out of which 1330 are citizens of Aarau.
 This also allows individuals to have a voice, who are not living in Aarau, but have a connection to the city, such as workers and broader family members.
 Children over 15 years old could also vote.
-Voting was digital, using the open-source Stanford Participatory Budgeting platform [43].
+Voting was digital, using the open-source Stanford Participatory Budgeting platform [Wellings2023].
 However, voters with low digital literacy were supported in virtual voting centers around the city.
-The voting platform was extended to support cumulative voting by distributing 10 points over at least 3 projects [30, 43].
+The voting platform was extended to support cumulative voting by distributing 10 points over at least 3 projects [SPBStadtidee2023, Wellings2023].
 Voters are securely verified in a privacy-preserving way by providing a unique (hash) code that was generated by the city council using the social security number of each citizen (AHV card).
 The same platform was used for project campaigning by using the pairwise project comparisons feature.
-The winning projects were calculated using the method of equal shares and the Add1U completion method implemented within the Pabutools (the ones of Świecie and Wieliczka used the Add1 completion method, while Add1U would yield 1 and 3 more winners) [37].
+The winning projects were calculated using the method of equal shares and the Add1U completion method implemented within the Pabutools (the ones of Świecie and Wieliczka used the Add1 completion method, while Add1U would yield 1 and 3 more winners) [pabutools].
 Voting outcomes with the utilitarian completion method were also calculated and assessed in the exit survey (Table LABEL:table:exit-voting_outcome).
 4.2 Evaluation of legitimacy
-To assess the legitimacy of ballot aggregation methods, we rely on well-established theory of political legitimacy [34] and social choice [44, 45].
-Among the three notions of input (voting participation), throughput (the voting process) and output (voting outcome) legitimacy [46, 47], we mainly focus in Section 2.2 on output legitimacy.
+To assess the legitimacy of ballot aggregation methods, we rely on well-established theory of political legitimacy [Weatherford1992] and social choice [Arrow2010, Brandt2016].
+Among the three notions of input (voting participation), throughput (the voting process) and output (voting outcome) legitimacy [Scharpf1999, Schmidt2013], we mainly focus in Section 2.3 on output legitimacy.
 However, we also assess the choice of ballot aggregation method before and after voting with equal shares as well as the effect of explaining equal shares, both capturing throughput legitimacy.
 As such, by legitimacy we refer to voters who remain satisfied and accept the voting outcome of a ballot aggregation method (explained or not) even if they are among the losers, i.e.
-the did not vote for a winning project [48].
+the did not vote for a winning project [Hanggli2024].
 By considering all 24 combinations of winner vs.
@@
 We use different satisfaction proxies based on questions of the exit survey in Table LABEL:table:entry-digital (OB.1).
-In particular, we assess whether voters are satisfied with the voting outcome of City Idea (OB.1.1), whether they accept this outcome (OB.1.2) and whether they feel the outcome accurately represents the will of Aarau citizens (OB.1.7).
+In particular, we assess whether voters are satisfied with the voting outcome of “City Idea” (OB.1.1), whether they accept this outcome (OB.1.2) and whether they feel the outcome accurately represents the will of Aarau citizens (OB.1.7).
 Based on 11-point likert-scale responses, we split the populations two quartiles, the first one for the satisfied voters and the second one for the unsatisfied ones.
@@
 4.3 Explaining shifts between ballot aggregation methods
-In this section, we present a causal analysis of how personal traits influence shifts in individual voter preferences for aggregation methods, based on the pre-voting survey, the actual voting process and the post-voting survey.
+In this section, we present a causal analysis of how personal traits influence shifts in individual voter preferences for ballot aggregation methods, based on the pre-voting survey, the actual voting process and the post-voting survey.
 Prediction modeling for causal explanations.
@@
 Prediction algorithms.
-Prediction models are constructed using tree-based gradient boosting algorithms [49] and neural networks with three hidden layers [50].
+Prediction models are constructed using tree-based gradient boosting algorithms [prokhorenkova2018catboost] and neural networks with three hidden layers [singh2019study].
 Different groups of personal human traits are assessed for the capacity to causally explain shifts between different ballot aggregation methods.
 The selection of these algorithms aligns well with how they handle categorical data (e.g.
-CatBoost [51] and neural networks [52] known to perform well even on smaller datasets).
+CatBoost [hancock2020catboost] and neural networks [wu2018development] known to perform well even on smaller datasets).
 Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the models are illustrated in Section LABEL:sec:shifts-methods.
@@
 Once predictions are made, we derive the relative importance of the personal traits via a two-step explainable framework.
-Initially, an ablation study [53] is performed to systematically evaluate the relative predictive contribution of each personal human trait.
-In the second step, we employ model-agnostic Shapley Additive Explanations (SHAP) [54] to assess the individual contribution of each trait to the prediction performance.
+Initially, an ablation study [mohseni2021multidisciplinary] is performed to systematically evaluate the relative predictive contribution of each personal human trait.
+In the second step, we employ model-agnostic Shapley Additive Explanations (SHAP) [nohara2019explanation] to assess the individual contribution of each trait to the prediction performance.
 More details regarding the performance of the explanatory framework can be found in Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-lossLABEL:tab:trait-survey-map and Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
@@
 FBH: Software, Validation, Data Curation.
-RHF: Conceptualization, Methodology, Data Collection, Investigation, Resources, Supervision, Writing - Review & Editing, Project administration, Funding acquisition DH: Conceptualization, Investigation, Resources, Writing - Input to Original Draft, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
+RHF: Conceptualization, Methodology, Data Collection, Investigation, Resources, Supervision, Writing - Review & Editing, Project lead and administration, Funding acquisition DH: Conceptualization, Investigation, Resources, Writing - Input to Original Draft, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
 Competing Interests
@@
 Acknowledgments
-This project received support by the SNF NRP77 ‘Digital Transformation’ project "Digital Democracy: Innovations in Decision-making Processes", #407740_187249 and the UKRI Future Leaders Fellowship (MR/W009560/1): ‘Digitally Assisted Collective Governance of Smart City Commons–ARTIO’.
-The authors would like to thank the representatives from the city of Aarau for supporting and running the "Stadtidee" [translated: "City Idea"] campaign.
-The authors would also like to thank Abhinav Sharma for supporting the team at University of Leeds to collect the voting data as well as Michael Buess from Demoscope for his support with the surveys.
+This project received support by the SNF NRP77 ‘Digital Transformation’ project “Digital Democracy: Innovations in Decision-making Processes”, #407740_187249 and the UKRI Future Leaders Fellowship (MR/W009560/1): “Digitally Assisted Collective Governance of Smart City Commons–ARTIO”.
+The authors would like to thank the representatives from the city of Aarau for supporting and running the “Stadtidee” [translated: "City Idea"] campaign.
+The authors would also like to thank Abhinav Sharma for supporting the team at University of Leeds to collect the voting data, Sajan Maharjan for his support with the Pabulib data, as well as Michael Buess from Demoscope for his support with the surveys.
 The authors are grateful to Lodewijk Gelauff and Ashish Goel for their invaluable support they provided us to extend the Stanford Participatory Budgeting platform.